### v3

# Efficient Nearest Neighbor Search Using Dynamic Programming

Abstract
When dealing with point clouds distributed on manifold surfaces in 3D space, or when the query point is far from the data, the efficiency of traditional nearest neighbor search algorithms (e.g., KD Tree and R Tree) may degrade. In extreme cases, the complexity of the query can approach $O(n)$. In this paper, we propose a novel dynamic programming technique that precomputes a Directed Acyclic Graph (DAG) to enable more efficient nearest neighbor queries for 2D manifold data. By leveraging this structure, only a small number of distance comparisons between point pairs are required to accurately identify the nearest neighbor. Extensive experimental results demonstrate that our method achieves query speeds that are 1x-10x faster than traditional methods. Moreover, our algorithm exhibits significant potential. It achieves query efficiency comparable to KD-trees on uniformly distributed point clouds. Additionally, our algorithm supports nearest neighbor queries among the first $k$ points. Coupled with our algorithm, a farthest point sampling algorithm with lower complexity can also be implemented. Furthermore, our method has the potential to support nearest neighbor queries with different types of primitives and distance metrics. We believe that the method proposed in this paper represents the most concise and straightforward exact nearest neighbor search algorithm currently available, and it will contribute significantly to advancements in the field.
Index Terms:
Nearest neighbor search, farthest point sampling, density peak clustering, delaunay, voronoi diragram.
I Introduction
Given a target point set $\mathcal{P}=\{p_{i}\}_{i=1}^{n}$ and a query point $q$, the point $p_{i}$ is considered the nearest neighbor in $\mathcal{P}$ to $q$ if $p_{i}$ satisfies
$\forall j\in[1,n],\|q-p_{i}\|\leq\|q-p_{j}\|$ (1)
[…]
As shown in the inset figure, we have two point clouds with distinct distributions: one uniformly distributed within a unit cube, and the other distributed across the surface of the Earth. Additionally, we generate two sets of query points—one within the bounding box of the input data, and the other within 8x the bounding box. It can be observed that the query performance of both KD Tree and R Tree significantly degrades under these conditions.
In this paper, we present a simple, effective, and novel nearest point query algorithm based on Voronoi / Delaunay. For 2D manifold data in 3D, our query speed is 1x-10x faster than KD Tree and R Tree. Even for uniformly distributed random point clouds, our query algorithm achieves speeds comparable to KD Tree. Our main contributions are as follows:
  • We introduced a concise nearest neighbor search algorithm for 2D manifold point cloud data that significantly improves query speed compared to traditional algorithms.
  • If the point cloud possesses an order, our method can facilitate the identification of the nearest point to $q$ among $\{p_{i}\}_{i=1}^{m}$, for any $m$, where $m<n$. This approach can be employed to reduce the time complexity of Density Peak Clustering [9], thus enhancing its applicability in large-scale scenarios.
  • Based on our method, we developed a farthest point sampling algorithm with an average complexity of $O(k\log n)$ in 3D, where $n$ represents the total number of points and $k$ denotes the number of samples.
Additionally, our algorithm exhibits significant potential. For instance, under various primitive types and distance metrics, as long as the incremental Voronoi or Delaunay construction algorithms for the primitives, as well as the distance query function from a point in space to the primitive, are provided, our nearest neighbor search framework can support nearest neighbor searches in these contexts. However, further development will be necessary in future work.
II Releated Work
II-A Exact Nearest Neighbor Search
For spatial data, the nearest neighbor problem is usually solved by building a spatial index [10]. These indices can broadly be divided into two categories: those that partition the space and those that partition the data.
Classic structures for space partitioning include KD Tree and quadtrees. A KD Tree [7] is a spatial partitioning structure that recursively divides the dataset along one of the k coordinate system axes, typically alternating between them; for instance, with k = 2, splits occur alternately on the x- and y-axes. This splitting process continues until a maximum depth is reached or until fewer than a specified number of points remain in a cell. KD Trees can be extended to include objects beyond points, such as triangles or line segments. Determining the optimal splitting plane can be challenging in these cases; However, for points (or disks) in two dimensions, we can simply sort the points along each dimension and split at the median. The construction of a KD Tree can be achieved in $O(n\log n)$ time, and it can find the nearest neighbors in $O(\log n)$ time. A octree is another spatial partitioning structure, but it specifically divides a 3D space into eight cubes or regions. This is typically done recursively, with each subdivision occurring when the data within a region meets certain criteria, such as exceeding a maximum number of points or objects. Each internal node in the octree corresponds to a cube region of space, which is further subdivided into eight equal-sized cubes, represented by its children. The construction of an octree can be completed in $O(n\log n)$ time, and it can find the nearest neighbors in $O(\log n)$ time, too.
The R Tree [8] is a classic structure for partitioning data. It is based on B-trees and shares a similar structure: both the number of objects in a leaf node and the number of children in an internal node are bounded by specified minimum and maximum limits. Additionally, all leaf nodes are at the same level. The R-tree can be built in $O(n\log n)$ time and allows for the identification of the nearest neighbors in $O(\log n)$ time.
Grids and linear search methods are also classic approaches for nearest-neighbor searches. The grid-based approach improves efficiency by overlaying a regular grid on the entire simulation area and tracking the agents within each cell. During a query, the search includes the cell containing the query points and the adjacent cells. The construction time for this method is $O(n+m)$, where $m$ represents the number of cells in the grid. The worst-case query time is $O(n)$, but practical performance greatly depends on the distribution of agents and the size of the cells. On the other hand, linear search is suitable when the number of query points is very small. This method finds the nearest points by calculating the distance from each target point to the query point.
II-B Voronoi Diagram & Delaunay
Suppose that we have a set of generators $\mathcal{S}=\{s_{i}\}_{i=1}^{n}$ in a given domain $\Omega$ that is equipped with a metric function $\mathcal{D}$, the Voronoi diagram involves partitioning $\Omega$ into regions such that the generator $s_{i}$ dominates a region
$\text{Cell}(s_{i})=\{x\in\Omega~{}\big{|}~{}\mathcal{D}(s_{i},x)\leq\mathcal{D}(s_{j},x),j\neq i\}.$ (2)
The definition of Voronoi diagrams may vary with domains, metrics, and generator types. The most prevalent version assumes that $\Omega$ represents Euclidean spaces, and the generators are exclusively points. As a fundamental tool, this has found widespread applications in computer graphics [3, 1, 11].
The Delaunay is the dual structure of the Voronoi diagram, representing the adjacency relationships between the Voronoi cells. Currently, Voronoi diagrams are typically computed by first calculating the Delaunay triangulation and then deriving the dual structure. The Watson algorithm [12] is a classic method for computing the Delaunay triangulation. It is based on the empty circle property and works by incrementally inserting points and flipping triangles that do not satisfy this property. Common tools for computing Delaunay include TetGen [13] and CGAL [14], which can achieve an average computational complexity of $O(n\log n)$, with a worst-case complexity of $O(n^{2})$.
Voronoi diagrams can also be applied to nearest neighbor searches. By definition, if the closest point to a query point $q$ in $\mathbf{P}$ is $p_{k}$, then $q$ lies within the Voronoi cell of $p_{k}$. This property makes Voronoi diagrams highly relevant for solving nearest neighbor problems. However, the challenge lies in efficiently identifying the specific Voronoi cell containing $q$, as there currently isn’t an efficient algorithm for this task. This limitation restricts the practical use of Voronoi diagrams in nearest neighbor searches.
II-C Farthest Point Sampling
Farthest point sampling (FPS) is a technique widely employed in computational geometry and computer graphics applications. It is designed to iteratively select points from a dataset such that each newly selected point is maximally distant from the previously chosen points. This method can be formally expressed as follows:
$p_{i+1}=\arg\max_{p\in P}\left(\min_{p_{j}\in S}d(p,p_{j})\right),$ (3)
where:
  • $P$ represents the set of points in the point cloud,
  • $S$ denotes the subset of points that have already been selected,
  • $d(p,p_{j})$ is the distance between points $p$ and $p_{j}$,
  • $p_{i+1}$ is the next point selected such that its minimum distance to any point in $S$ is maximized.
This sampling strategy ensures that the selected points are well-distributed, making it effective for approximating shapes, surfaces, or distributions. FPS is particularly useful in applications requiring a representative subset of data points, such as mesh generation, surface simplification, or cluster initialization.
However, a notable drawback of FPS is its computational cost. The complexity of the method is $O(kn)$, where $n$ is the total number of points in the dataset, and $k$ denotes the number of points to be sampled. This complexity can become prohibitive for large datasets or when a high number of samples is needed, making FPS less efficient in such cases.
[…]
Figure 1: (a) The Voronoi diagram of $\{p_{i}\}_{i=1}^{20}$, where the query point $q$ (colored in red) belongs to the Voronoi cell colored in yellow. (b) If the addition of the three sites (whose cells are colored in green) does not change the yellow cell, $\Phi_{m}(q)$ remains unchanged when $m$ is increased from 20 to 23. (b) When the 24-th site $p_{24}$ is added (see the cyan cell), the yellow cell diminishes; yet $\Phi_{24}(q)=\Phi_{23}(q)$ if $q$ is nearer to the old site. (d) $\Phi_{24}(q)=p_{24}$ if $q$ is nearer to the newly added site $p_{24}$.
III Methodology
In this section, we introduce the key concept underlying the algorithms discussed in this paper, beginning with an exploration of linear search.
We assume that the given points are numbered from 1 to $n$. Given a query point $q$, we use $\Phi_{m}(q)$ to denote the closest point in $\{p_{i}\}_{i=1}^{m}$, where $m\leq n$. This leads to the following state transfer equation:
$\Phi_{m}(q)=\begin{cases}p_{1}&\text{if }m=1,\\ \Phi_{m-1}(q)&\text{if }m\neq 1\ \&\ \|q-\Phi_{m-1}(q)\|\leq\|q-p_{m}\|,\\ p_{m}&\text{if }m\neq 1\ \&\ \|q-\Phi_{m-1}(q)\|>\|q-p_{m}\|,\end{cases}$ (4)
which can be understood as a dynamic programming problem, where $\Phi_{n}(q)$ is the actual nearest point to be extracted. Obviously, a naïve implementation of the query operation costs $O(n)$ time. In the following, we consider how to boost this implementation.
Recall that Voronoi diagrams [15] precisely characterize how the space of interest is divided by the given data points. A point $q$ is nearest to $p_{i}$ if and only if $q$ is located in $p_{i}$’s cell. However, Voronoi diagrams are not well-suited for NNS due to their lack of hierarchical representation. Despite this, during the incremental construction process of Voronoi Diagram, the area dominated by a site either remains unchanged or shrinks, encoding the history of how the proximity structure changes when a new point is added. Let the input point cloud be $\mathcal{P}=\{p_{i}\}_{i=1}^{n}$. We illustrate our construction process by considering any moment of constructing the Voronoi diagram through incremental point insertion, such as $\mathcal{V}_{m}$, which is constructed from the set $\{p_{i}\}_{i=1}^{m}$, where $m\leq n$.
Figure 1(a) shows the Voronoi diagram $\mathcal{V}_{20}$ determined by 20 sites, with one of the cells highlighted in yellow. By adding three additional sites (whose cells are highlighted in green), we obtain a new Voronoi diagram $\mathcal{V}_{23}$; see Figure 1(b). It can be seen that the addition of the three green sites does not change the yellow cell. Therefore, if $q$ is located in the yellow cell, then $\Phi_{23}(q)=\Phi_{20}(q)$, implying that the nearest site to $q$ remains unchanged even when the three green sites are added. Only when a newly added site changes the yellow cell may the nearest site differ. Specifically, $\Phi_{24}(q)=\Phi_{23}(q)$ if $q$ is nearer to the old site (see Figure 1(c)), and $\Phi_{24}(q)=p_{24}$ if $q$ is nearer to the new site $p_{24}$ (see Figure 1(d)). To summarize, by using $p_{m+1}\nvdash\text{Cell}(\Phi_{m}(q);\mathcal{V}_{m})$ to represent that the addition of $p_{m+1}$ does not diminish the cell of $\Phi_{m}(q)$ in $\mathcal{V}_{m}$, we have:
$p_{m+1}\nvdash\text{Cell}(\Phi_{m}(q);\mathcal{V}_{m})\quad\mapsto\quad\Phi_{m+1}(q)=\Phi_{m}(q),$ (5)
or
$p_{m+j}\nvdash\text{Cell}$ $(\Phi_{m}(q);\mathcal{V}_{m+j-1}),\ \forall 1\leq j\leq k\quad$
[…]
$\Phi_{m+k}(q)$ $=\Phi_{m+k-1}(q)=\cdots=\Phi_{m}(q).$
IV Algorithm
IV-A Query Table Construction
Based on Section III, we can design a data structure called the Query Table, denoted as =$\mathcal{L}=\{\mathcal{L}_{i}\}_{i=1}^{n}$, to store the necessary information for the query phase. Specifically, we assign each point $p_{i}$ a corresponding vector $\mathcal{L}_{i}$, called Query List. During the incremental construction of the Voronoi diagram, when $p_{j}$ is inserted, if the condition $p_{j}\vdash\text{Cell}(p_{i};\mathcal{V}_{j-1})$ is satisfied, we push $p_{j}$ to the end of $\mathcal{L}_{i}$.
Query Table $\mathcal{L}$ is constructed synchronously as the Voronoi diagram is incrementally built. Initially, we have only one site $p_{1}$, leaving its Query List empty. Suppose that the addition of $p_{m+1}$ changes $p_{i}$’s cell in $\mathcal{V}_{m}$. We append $p_{m+1}$ to $p_{i}$’s Query List; see Figure 2. The Query Table $\mathcal{L}$ is completed once all the sites are inserted. It is important to note that the order of elements stored in each $\mathcal{L}_{i}$ are consistent with the insertion order.
[…]
Figure 2: We consider an incremental construction process of the Voronoi diagram of $\{p_{i}\}_{i=1}^{n}$. During the process, we append $p_{j}$ to $p_{i}$’s Query List if the addition of $p_{j}$ changes $p_{i}$’s cell. The rightmost figure shows the query process.
Since the Delaunay is the dual of the Voronoi diagram and reflects the adjacency relationships between Voronoi cells, to check whether $p_{m}\vdash\text{Cell}(p_{i};\mathcal{V}_{m-1})$, this can be further converted into checking whether $p_{m}$ and $p_{i}$ are adjacent in the Delaunay constructed from $\{p_{i}\}_{i=1}^{m}$, which is dual to $\mathcal{V}_{m}$. Algorithm 1 provides the pseudocode for the corresponding construction process.
Input: Point cloud $\mathcal{P}=\{p_{i}\}_{i=1}^{n}$.
Output: Query Table $\{\mathcal{L}_{i}\}_{i=1}^{n}$.
Initialize the Query Table =$\mathcal{L}=\{\mathcal{L}_{i}\}_{i=1}^{n}$.
Insert $p_{1}$, initialize the Delaunay structure $\mathbf{D}$.
foreach $i$ in $[2,n]$ do
     Initialize the set $\omega$ as empty.
     Insert $p_{i}$ and update $\mathbf{D}$.
     Extract all points adjacent to $p_{i}$, store them in $\omega$.
     foreach $p_{x}$ in $\omega$ do
           Push $p_{i}$ to the back of $\mathcal{L}_{x}$.
           end foreach
           end foreach
Algorithm 1 Query Table Construction
IV-B Nearest Neighbor Query
In the query phase, we initialize the nearest site as $p_{1}$, which implies $\Phi_{1}(q)=p_{1}$. Let $p_{2}$ be the first site in $p_{1}$’s Query List $\mathcal{L}_{1}$, and we visit $\mathcal{L}_{1}$ in order. If $p_{2}$ provides a smaller distance than $p_{1}$, we switch to $p_{2}$’s Query List $\mathcal{L}_{2}$. Otherwise, we move to the next site in $\mathcal{L}_{1}$ to see if it can provide a smaller distance. See the rightmost figure of Figure 2 for how the query algorithm operates. See Algorithm 2 for the pseudo-code of the query algorithm.
Input: Query Table $\{\mathcal{L}_{i}\}_{i=1}^{n}$ and query point $q$.
Output: The closest point to $q$.
getNearestNeighbor($p_{I}=p_{1}$)
foreach $p$ in $\mathcal{L}_{I}$ do
      if $\|q-p\|<\|q-p_{I}\|$ then
           return getNearestNeighbor($p$)
           end if
           end foreach
return $p_{I}$
Algorithm 2 Nearest Neighbor Search
IV-C Farthest Point Sampling
Sections IV-A and  IV-B describe the algorithm for constructing the Query Table and performing nearest neighbor searches. The algorithm assumes a specific order of the point cloud and incrementally constructs the Voronoi / Delaunay based on this order. When the point cloud is unordered or there is no need to compute the nearest point to $q$ among the first $n$ points, we can sort the points in a specific manner to enhance the efficiency of the query phase.
Intuitively, the more uniformly the inserted point cloud is distributed in space, the faster the query speed. This is because it facilitates a transition from large to small spatial jumps during the query process. Farthest point sampling, a commonly used downsampling method, aims to distribute the sampled points as uniformly as possible across the point cloud. Drawing inspiration from this, we set the number of sampled points equal to the total number of points in the cloud and use the sampling order to incrementally construct the Voronoi diagram. Notably, by integrating this approach into our construction process, we developed a farthest point sampling strategy with reduced complexity, leading to faster sampling times.
An important step in the farthest point sampling process involves calculating the distance between the newly sampled point and all remaining points, followed by updating the distance information. However, in most cases, the distances for the majority of the remaining points do not require updating, leading to a significant waste of computational resources. By integrating this process with our construction method, we can optimize the calculations by updating the distances for only a small subset of remaining points that are likely to change. This approach improves the algorithm’s efficiency.
As an example, consider a state from the incremental construction of the Voronoi diagram, as shown in Figure 3(a). The black points represent the points that have already been inserted, while the gray points denote the remaining points waiting to be inserted. When a new point is inserted based on the farthest point sampling strategy, as illustrated in Figure 3(b), the red point represents the newly sampled point from the gray points, and the green area corresponds to its Voronoi cell. A key observation is that only the distances of the remaining points located within the cells adjacent to the green cell (specifically, those within these cells prior to the insertion of the red point) need to be updated. Since the white Voronoi cell does not change during the insertion of new points, this implies that the nearest neighbors of the uninserted points within it remain unchanged. This significantly reduces the computational cost of sampling. The remaining task is to dynamically track the Voronoi cells that contain the remaining points as the Voronoi diagram is incrementally constructed. This process is straightforward.
[…]
Figure 3: Coupling of Farthest Point sorting with incremental Voronoi Construction. Left: Uninserted points (gray) are distributed within the Voronoi cells of the inserted points (black). Right: When a new point is inserted, only the distance of the uninserted points within the green and yellow cells may need to be updated.
IV-D Complexity Analysis
In this section, we will analyze the complexity of the farthest point sampling as a standalone component, rather than as part of the construction phase.
Time Complexity of Query Table Construction. The complexity of the Query Table Construction is the same as that of the incremental Delaunay construction. On average, the time complexity for constructing a Delaunay in 3D is $O(n\log n)$, while the worst-case time complexity is $O(n^{2})$. However, this worst-case scenario is seldom encountered in typical construction processes.
Space Complexity of the Query Table. On average, the number of neighbors for a Voronoi cell in 3D is a constant, denoted as $c$. This implies that when a new point is inserted, it is adjacent to $c$ cells on average, requiring the point to be appended to the end of $c$ Query Lists. Therefore, the average space complexity of the Query Table is
$O(\sum_{i=2}^{n}c)=c*O(n)=O(n),$ (7)
where $n$ is the number of points.
Time Complexity of Nearest Neighbor Query. When querying the nearest neighbor for the point $q$, the complexity of the query phase depends on the number of times the cell containing $q$ (which may change during the query process) shrinks during the incremental construction, as a comparison is needed each time the Voronoi cell containing $q$ shrinks. Simply put, assuming $\Phi_{m}(q)$ is known and a new point $p_{m+1}$ is inserted, the cell of $p_{m+1}$ typically adjacent to $c$ cells on average. Therefore, the probability that $\text{Cell}(\Phi_{m}(q);\mathcal{V}_{m})$ shrinks is approximate $c/m$, indicating the possibility of needing a comparison. Thus, the average time complexity of the nearest neighbor query phase is
$O(\sum_{i=1}^{n-1}\frac{c}{i})=c*O(\log n)=O(\log n).$ (8)
Time Complexity of Farthest Point Sampling. When we have already inserted $k$ points to construct the Voronoi diagram, each cell, on average, contains ${(n-k)}/{k}$ remaining points. Given the distances from all remaining points to the nearest sampled points are known, selecting the next point for sampling has a complexity of $O(\log(n-k))$. Knowing that the newly inserted point is adjacent to $c$ cells, the number of remaining points that potentially need their distances updated is $c\times(n-k)/k$.
Thus, the additional complexity caused by the farthest point sampling process is:
$O(\sum_{k=1}^{n-1}\left[\log(n-k)+\left(c*\frac{n-k}{k}\right)\right])$ (9)
$=O(\sum_{k=1}^{n-1}\log(k)+c*n*\sum_{k=1}^{n-1}\frac{1}{k}-(n-1)*c)$ (10)
$=O(n\log n)$ (11)
V Results
Hardware environment. Our algorithm was implemented in C++ on a platform featuring a 3.4 GHz AMD Ryzen 9 5950X 16-Core CPU, 64GB of RAM, and running the Windows 11 operating system.
Experimental parameters. We utilized CGAL [14] for Delaunay computations. In our experiments, we used random sorting and farthest point sorting as insertion order. We measured and recorded both the construction performance and query performance for each method. The space for randomly generating query points was confined within the bounding box with 8x the volume (with edge lengths twice that of the original bounding box) of the input point cloud. The number of query points generated was set at 1000K.
Relevant approaches for comparison. Currently, there are numerous algorithms [16, 17, 18, 19] available for nearest neighbor search. In this paper, we selected two of the fastest and most classic methods for comparison.
  1. 1.
    KD Tree: We utilized the implementation from the nanoflann library [20], setting the maximum number of leaf nodes to 10. According to tests referenced in [21], this implementation is currently the fastest for nearest neighbor searches.
  2. 2.
    R Tree: We used the implementation from Boost Library [22].
In all the experiments below, Ours1 refers to our method using a random order for incremental Delaunay construction, while Ours2 indicates that the farthest point sorting strategy was applied. All the point cloud data used in the experiment has been included in the supplementary data.
V-A Query Table
Length of the query list for a vertex
Recall that the Query Table maintains a list of points encroached upon during the incremental construction of the Delaunay for each vertex in $V$. It is evident that the Voronoi cells of points inserted earlier often shrink more times, resulting in longer Query Lists. Therefore, to investigate how the length of the Query List varies with different insertion indexes, we conducted a study using the $100$K-vertex dragon point cloud as an example. As shown in Figure 4, we have plotted the length of the query list for each point with respect to its insertion order. Although the impact of random sorting and farthest point sorting strategies on the length of the Query List is not readily apparent from the figure, we calculated the average lengths of the Query Lists to be 15.684 and 14.165, with standard deviations of 17.705 and 16.491, respectively. The farthest point sorting resulted in a shorter and more uniform distribution of the Query List. Similar results were obtained across multiple data. This characteristic enables the query table generated through farthest point sorting to facilitate faster nearest-neighbor searches.
[…]
Figure 4: The length of the Query List for different index points. Top: random sorting. Bottom: farthest point sorting.
Number of tested vertices
As shown in Algorithm 2, the nearest neighbor search problem is transformed into the task of calculating distances between points.
The number of distance comparisons between points during the query process partially reflects the efficiency of the nearest neighbor search. We tested several classical models and recorded the number of comparisons, along with the mean and variance of the Query Lists, as shown in Table I. It can be observed that, after utilizing the farthest point sorting method, both the average length (Avg) and variance (Var) of the Query Lists decreased, accompanied by a reduction in the average number of comparisons during the nearest neighbor search process.
TABLE I: The average (Avg) and variance (Var) of the length of the query list and the number of comparisons required per query on classical models.
Camel Bunny Dragon Kitten Armadillo Lucy Sponza
Vertices 28934 72911 100313 291023 726367 1018219 1313504
$\text{Ours}^{1}$ Avg 15.913 15.805 15.684 15.397 15.832 16.034 14.631
Var 17.911 17.264 17.705 18.834 17.631 17.968 20.431
Tested 171.783 166.726 209.516 182.719 234.742 276.549 418.285
$\text{Ours}^{2}$ Avg 14.414 14.345 14.165 13.964 14.370 14.478 13.067
Var 16.783 15.906 16.491 17.126 16.172 16.584 19.681
Tested 144.754 135.831 161.822 155.806 197.042 235.085 358.413
V-B Preprocessing Cost
When a random sorting method is adopted, the preprocessing time of the current algorithm is equivalent to that of incremental Delaunay construction. However, at this stage, our preprocessing speed is slower than that of existing Delaunay construction libraries. This is primarily due to two factors. First, our method currently supports only serial execution. Second, certain algorithms, such as TetGen, reorder points before incremental Delaunay construction to enhance overall speed. However, this insertion order is not favorable for our queries. It is worth noting, however, that our method has the potential to accelerate preprocessing when integrated with Delaunay construction. In the current implementation, we present the preprocessing times of different algorithms across several classic models in Table II. Additionally, Figure 5 shows the preprocessing time with respect to the resolution of the Dragon data, comparing our method to other methods. It can be observed that at this stage, the preprocessing time of our method exceeds that of existing nearest neighbor search algorithms.
TABLE II: Preprocessing time ($s$) on classical models.
Camel Bunny Dragon Kitten Armadillo Lucy Sponza
Vertices 28934 72911 100313 291023 726367 1018219 1313504
KD-tree 0.004 0.011 0.016 0.051 0.149 0.215 0.281
R-tree 0.005 0.014 0.022 0.062 0.170 0.243 0.311
Ours1 1.097 3.205 4.692 14.46 41.58 62.13 72.31
Ours2 1.876 5.818 8.473 35.64 96.11 263.3 307.6
[…]
Figure 5: The preprocessing cost of KD Tree, R Tree and Ours on the Dragon data with varying resolutions. Top: preprocessing time. Bottom: the comparison about the preprocessing cost. Dashed line indicates where the ratio equals 1.
V-C Query Performance
Using the Dragon dataset as an example, we present the query cost of our method compared to other approaches with respect to resolution in Figure 6, demonstrating the significant advantage of our method in this scenario. In Table III, we provide more detailed examples for several classical models.
[…]
Figure 6: Average query time on the Dragon model with varying resolutions. Top: Average query time ($\mu$s) among KD tree, R tree and ours. Bottom: The comparison about the average time different methods. Dashed line indicates where the ratio equals 1.
TABLE III: Average query time ($\mu s$) of KD Tree, R Tree, and Ours on classical models. The highest and second-fastest query times are highlighted in bold and underlined, respectively.
Camel Bunny Dragon Kitten Armadillo Lucy Sponza
Vertices 28934 72911 100313 291023 726367 1018219 1313504
KD Tree 1.639 3.795 2.475 8.475 12.81 9.796 4.407
R Tree 1.623 2.739 2.034 5.359 5.241 4.518 1.851
Ours1 0.423 0.559 0.540 0.957 1.132 1.293 2.146
Ours2 0.344 0.482 0.388 0.752 0.834 1.070 1.694
V-D Tests on ABC & Thingi10K
Thingi10K [23] and ABC [24] are two large-scale 3D datasets that contain diverse models. To conduct a comprehensive comparison among KD Tree, R Tree, and Ours, we ran them on the dataset to compare both the preprocessing cost and the query cost. The query time are shown in Figures 7 and Figure 8. And the preprocessing time can be found in Figure 13 and Figure 14 in Appendix.
[…]
Figure 7: Comparison about query performance on ABC dataset. Top: the average timing cost per query ($\mu$s) for KD Tree, R Tree and Ours. The y-axis is scaled logarithmically. Bottom: the comparison about the query cost among KD Tree, R Tree and Ours. Dashed line indicates where the ratio equals 1.
[…]
Figure 8: Comparison about query performance on Thingi10K data. Top: the average timing cost per query ($\mu$s) for KD Tree, R Tree and Ours. The y-axis is scaled logarithmically. Bottom: the comparison about the query cost among KD Tree, R Tree and Ours. Dashed line indicates where the ratio equals 1.
V-E Other Types of Data
Uniform Point Cloud.
To fully demonstrate the efficiency of the algorithm, we compared KD Tree, R Tree, and Ours on uniform point clouds. As shown in Figure 9, it can be observed that even on completely uniform point clouds, our method remains comparable in query efficiency to the fastest current implementation of KD Tree and R Tree.
[…]
Figure 9: Average query time on the uniform point cloud with varying resolutions. Top: Average query time per query ($\mu$s) among KD Tree, R Tree and Ours. Bottom: The comparison of the average time of different methods. Dashed line indicates where the ratio equals 1.
Earth City.
Additionally, we tested the efficiency of the KD Tree, R Tree, and our method on an Earth data, distributed over a sphere, with the results shown in Figure 10. At this distribution, both KD Tree and R Tree exhibit a significant drop in efficiency. However, our method remains extremely fast, achieving up to 40x the query speed of KD Tree.
[…]
Figure 10: Average query time on the earth data with varying resolutions. Top: Average query time among KD tree, R tree and ours. Bottom: The comparison about the average time different methods. Dashed line indicates where the ratio equals 1.
V-F Near and far query points.
Additionally, we observed that the location of query points at different positions can impact the efficiency of the query. To further investigate this, we conducted tests on uniformly distributed point clouds, point clouds representing Earth cities, and the Dragon model, with the number of points in each point cloud controlled at 1, [^0] K. The results are presented in Figure 11. The horizontal axis represents the ratio of the length of the generation space for the query points to the length of the bounding box of the target points along any coordinate axis. It can be observed that even in a completely uniform point cloud distribution, where the query point generation space matches the bounding box of the target points, our query speed remains comparable to that of the KD Tree. In other scenarios, our method demonstrates a significant advantage in query speed.
[…]
Figure 11: Average query time with varying sizes of the sampling box (arbitrary axis ratio, with volume ratio calculated as the cube of the axis ratio). Top: Uniform point cloud. Middle: Earth city data. Bottom: Dragon point cloud.
VI Desirable Properties
Beyond its speed advantages, we believe the current method holds significant potential, and future high-quality research could further expand the algorithm’s capabilities. In the following section, we will analyze several key aspects.
VI-A Nearest Neighbor Among the First $n$ Points
One notable property of our algorithm is that, given any query point $q$ and integer $k$, we can efficiently identify the nearest point among $\{p_{i}\}_{i=1}^{k},k\leq n$. Specifically, the order of the given points is used to incrementally construct the Delaunay. As discussed in Section III, during the nearest-neighbor search process, when querying the nearest neighbor, we only need to compute $\Phi_{n}(q)$. However, in practice, $\forall~{}i\in[1,n]$, $\Phi_{i}(q)$ is already incorporated into our query process. When a specific $k$ is provided, as shown in Algorithm 2, our query process can terminate early. To the best of our knowledge, no existing algorithm is capable of achieving this.
The Density Peak Clustering algorithm [9] is a well-known method for clustering. One critical step in this algorithm is the calculation of $\sigma_{i}$, which is defined as:
$\delta_{i}=\min_{j:\rho_{j}>\rho_{i}}(d_{ij}),$ (12)
$\rho_{i}$ is defined as the density of point $p_{i}$, and $d_{ij}$ represents the distance between points $p_{i}$ and $p_{j}$. Generally, computing $\{\delta_{i}\}_{i=1}^{n}$ requires calculating the distance between every pair of points, which has a complexity of $O(n^{2})$. However, for 3D points, using our algorithm, this problem can be solved in $O(n\log n)$ time.
VI-B Farthest Point Sampling.
Farthest Point Sampling (FPS) is a widely used sampling algorithm due to its ability to ensure uniform sampling of points, making it broadly applicable. For instance, in the PointNet++ [25] framework for 3D point cloud deep learning, FPS is employed to sample and cluster points as part of the receptive field. The complexity of sampling $k$ points from $n$ points is $O(kn)$. Currently, there are approximate farthest point sampling algorithms designed for specific types of point clouds [26], but as far as we know, no breakthroughs have been made in reducing the complexity of exact farthest point sampling. However, based on our algorithm, we achieve a complexity of $O(k\log n)$ for sampling $k$ points from $n$ points in 3D space.
VI-C Various Primitives & Distance Metrics
Many existing nearest neighbor search algorithms are designed for specific types of primitives and distance metrics, which limits their generalizability across diverse scenarios. In contrast, our method, as described in Sections IIIIV-A, and IV-B, requires only the incremental construction of the Voronoi diagram for the primitives and a distance query function from any point in space to the primitives, making it straightforward to integrate into our nearest neighbor query framework. For instance, by supplying only the algorithm for incrementally constructing the Voronoi diagram of triangles under the Manhattan distance, along with a function to query the Manhattan distance between a spatial point and a triangle, our framework can efficiently determine the nearest primitive to a given point under the Manhattan distance. We believe this is a significant advantage of our algorithm. Of course, we acknowledge that the incremental construction of Voronoi diagrams for various primitives is quite challenging; however, as outlined below, we also provide alternative options.
VI-D Fast Preprocessing & Slow Query
We acknowledge that the preprocessing efficiency of our algorithm is relatively low, especially in high-dimensional spaces. Furthermore, as previously mentioned, the incremental construction of Voronoi diagrams for various primitives under different distance metrics presents significant challenges. Therefore, we propose a new direction. As discussed in Section IV-A, when constructing the Query Table, we need to identify the set $\omega_{m}$ upon the insertion of $p_{m}$, as shown in Algorithm 1. The set is defined as follows:
$\omega_{m}=\{p_{i}~{}|~{}p_{m}\vdash\text{Cell}(p_{i};\mathcal{V}_{m-1})\}.$ (13)
In fact, when dealing with complex high-dimensional or intricate scenarios, computing $\omega_{m}$ becomes extremely challenging. Therefore, we compute $\omega_{m}^{*}$ instead, which only needs to satisfy $\omega\subseteq\omega_{m}^{*}$, while still ensuring accurate nearest neighbor results. This approach offers a direction for performing fast nearest neighbor searches in high-dimensional spaces and scenarios involving different primitives and distance metrics. As a result, the computed set may contain redundancies, leading to a slight reduction in the speed of the nearest neighbor search. Additionally, we believe that even if $\omega_{m}$ does not belong to $\omega_{m}^{*}$, it could provide valuable insights for developing approximate nearest neighbor search algorithms.
VI-E Fast Query & Slow Preprocessing
Another encouraging finding is that the number of comparisons in our query process, as shown in Table I, can be further reduced, suggesting potential for additional algorithmic acceleration. Specifically, as previously outlined in our algorithm, when we focus on the time point $\Phi_{k}(q)=p_{k}$ during the query process, the current approach only utilizes the information that $q\in\text{Cell}(\Phi_{k}(q);\mathcal{V}_{k})$. However, we can infer that $q\in\cap_{\Phi_{i}(q)=p_{i},1<i\leq k}\text{Cell}(\Phi_{i-1}(q);\mathcal{V}_{i-1})~{}\cap~{}\text{Cell}(\Phi_{k}(q);\mathcal{V}_{k})$ from the query process.
As illustrated in Figure 12, (a) and (b) depict a simple incremental construction process, with the red dot representing the query point $q$. In this example, we have only utilized the information shown in (c), namely that $q\in\text{Cell}(p_{3};\mathcal{V}_{3})$. As shown in brown in (c). However, based on the query process, we can infer that $q\in\text{Cell}(p_{1};\mathcal{V}_{2})~{}\cap~{}\text{Cell}(p_{3};\mathcal{V}_{3})$. As shwon in (d). This enables us to design a more rigorous and precise occupancy evaluation. Our method currently does not leverage this information during the query process, even though it can be obtained through complex preprocessing, which would, of course, involve substantial preprocessing overhead.
[…]
Figure 12: (a) and (d): A simple incremental insertion process, with the red dot representing the query point $q$. (c) Based on the query process of (a)(b), we know the information $q\in\text{Cell}(p_{3};\mathcal{V}_{3})$, as shown in brown. (d) In fact, the complete information we can obtain is $q\in\text{Cell}(p_{1};\mathcal{V}_{2})~{}\cap~{}\text{Cell}(p_{3};\mathcal{V}_{3})$, as shown in brown.
However, we also recognize that utilizing all of the aforementioned information would result in significantly increased preprocessing complexity. To balance this, we implemented a simplified approach that compromises between the algorithm described in this paper and the information mentioned in this section. In brief, in addition to $q\in\text{Cell}(p_{i};\mathcal{V}_{i})$, we incorporated $q\in\text{Cell}(\Phi_{i-1};\mathcal{V}_{i-1})~{}\cap~{}\text{Cell}(p_{i};\mathcal{V}_{i})$. This region is used as the fundamental unit for occupancy evaluation, rather than a complete Voronoi cell.
We were pleasantly surprised to find that, despite implementing only a simplified version, we achieved a significant reduction in the number of comparisons compared to Table I. As shown in Table IV, $Ours^{1}_{*}$ and $Ours^{2}_{*}$ represent the number of comparisons without and with farthest point sampling, respectively, under the strategy outlined in this section. Remarkably, only around 100 pairwise distance comparisons are required to find the exact nearest neighbor for a query point among 1 million points. However, we must acknowledge that while the number of comparisons has decreased, the actual query time has increased. We speculate that this may be due to increased memory usage by the data structure during the query phase, which likely results in more cache misses, thereby reducing efficiency. Nevertheless, we believe this is a highly meaningful direction for further exploration.
TABLE IV: The number of pairwise comparisons during the query process using different construction strategies.
Camel Bunny Dragon Kitten Armadillo Lucy Sponza
Vertices 28934 72911 100313 291023 726367 1018219 1313504
$\text{Ours}^{1}$ 171.783 166.726 209.516 182.719 234.742 276.549 418.285
$\text{Ours}^{2}$ 144.754 135.831 161.822 155.806 197.042 235.085 358.413
$\text{Ours}^{1}_{*}$ 93.218 97.49 108.545 107.708 127.760 134.123 124.615
$\text{Ours}^{2}_{*}$ 75.453 74.055 88.089 82.812 94.953 106.269 124.547
VII Limitations and Future Work
The main drawback of our algorithm at present is its preprocessing cost, which limits its application in some scenarios. However, fortunately, our method supports nearest neighbor queries during the construction process, which may assist in the tetrahedron location problem during incremental Delaunay construction, offering a potential avenue for accelerating preprocessing. Therefore, in future work, we will focus on improving the preprocessing speed. Additionally, our method currently only supports nearest neighbor search; we aim to extend the current algorithm to support k-nearest neighbor queries in the future. Furthermore, all the properties mentioned in Section VI are goals we aim to address in our future work.
VIII Conclusion
In this paper, we observe that when dealing with point clouds distributed over 2D manifolds in 3D space, or when query points are distant from the target points, the efficiency of traditional nearest neighbor search algorithms decreases. In certain special cases, the complexity can degrade to $O(n)$. Inspired by the incremental Delaunay construction process, we propose a nearest neighbor search algorithm specifically for 2D maniflod point cloud. Our method achieves 1-10 times the speed of KD Tree in nearest neighbor searches. We validated the efficiency advantages of our approach through extensive experiments.
Additionally, our algorithm demonstrates significant potential. For instance, it supports querying the closest point to $q$ among any of the first $k$ points. This capability facilitates the application of peak clustering algorithms to large-scale data scenarios. Moreover, we have implemented a farthest point sampling algorithm with $O(k\log n)$ complexity in 3D space. Our mathod can efficiently identify the point closest to the query point from a set of 1 million points by performing about 100 pairwise distance comparisons.
We believe that the algorithm presented in this paper holds significant potential for application and will make a substantial contribution to the community.
[…]
Figure 13: Comparison about preprocessing cost on ABC dataset. Top: the time cost of preprocessing for KD Tree, R Tree and ours. Bottom: the comparison about the preprocessing cost among KD Tree, R Tree and ours.
[…]
Figure 14: Comparison about preprocessing cost on Thingi10K dataset. Top: the time cost of preprocessing for KD Tree, R Tree and ours. Bottom: the comparison about the preprocessing cost among KD Tree, R Tree and ours.

### v4

@@
 Abstract
-When dealing with point clouds distributed on manifold surfaces in 3D space, or when the query point is far from the data, the efficiency of traditional nearest neighbor search algorithms (e.g., KD Tree and R Tree) may degrade.
-In extreme cases, the complexity of the query can approach $O(n)$.
-In this paper, we propose a novel dynamic programming technique that precomputes a Directed Acyclic Graph (DAG) to enable more efficient nearest neighbor queries for 2D manifold data.
-By leveraging this structure, only a small number of distance comparisons between point pairs are required to accurately identify the nearest neighbor.
-Extensive experimental results demonstrate that our method achieves query speeds that are 1x-10x faster than traditional methods.
-Moreover, our algorithm exhibits significant potential.
-It achieves query efficiency comparable to KD-trees on uniformly distributed point clouds.
-Additionally, our algorithm supports nearest neighbor queries among the first $k$ points.
-Coupled with our algorithm, a farthest point sampling algorithm with lower complexity can also be implemented.
-Furthermore, our method has the potential to support nearest neighbor queries with different types of primitives and distance metrics.
-We believe that the method proposed in this paper represents the most concise and straightforward exact nearest neighbor search algorithm currently available, and it will contribute significantly to advancements in the field.
+Given a collection of points in $\mathbb{R}^{3}$, KD-Tree and R-Tree are well-known nearest neighbor search (NNS) algorithms that rely on space partitioning and spatial indexing techniques.
+However, when the query point is far from the data points or the data points inherently represent a 2-manifold surface, their query performance may degrade.
+To address this, we propose a novel dynamic programming technique that precomputes a Directed Acyclic Graph (DAG) to encode the proximity structure between data points.
+More specifically, the DAG captures how the proximity structure evolves during the incremental construction of the Voronoi diagram of the data points.
+Experimental results demonstrate that our method achieves a 1x-10x speedup.
+Additionally, our algorithm offers several valuable features.
+For instance, it naturally supports an $O(k\log n)$ algorithm for farthest point sampling, where $k$ is the desired number of sample points.
+Moreover, density peak clustering, which involves finding the nearest point among the top K points, is typically considered to have a time complexity of $O(n^{2})$.
+With our algorithm, this can be reduced to $O(n\log n)$.
+We believe this work will inspire further research on the NNS problem.
 Index Terms:
@@
 Given a target point set $\mathcal{P}=\{p_{i}\}_{i=1}^{n}$ and a query point $q$, the point $p_{i}$ is considered the nearest neighbor in $\mathcal{P}$ to $q$ if $p_{i}$ satisfies
-$\forall j\in[1,n],\|q-p_{i}\|\leq\|q-p_{j}\|$ (1)
+$\forall j\in[1,n],\|q-p_{i}\|\leq\|q-p_{j}\|,$ (1)
 […]
+As shown in the inset figure, we consider two point clouds with distinct distributions: one uniformly distributed within a unit cube and the other across the surface of the Earth.
+Additionally, we generate two sets of query points—one within the bounding box of the input data and the other within 8 times the bounding box.
+It is evident that the query performance of both KD Trees and R Trees deteriorates significantly under these conditions.
 […]
-As shown in the inset figure, we have two point clouds with distinct distributions: one uniformly distributed within a unit cube, and the other distributed across the surface of the Earth.
-Additionally, we generate two sets of query points—one within the bounding box of the input data, and the other within 8x the bounding box.
-It can be observed that the query performance of both KD Tree and R Tree significantly degrades under these conditions.
 In this paper, we present a simple, effective, and novel nearest point query algorithm based on Voronoi / Delaunay.
@@
   • We introduced a concise nearest neighbor search algorithm for 2D manifold point cloud data that significantly improves query speed compared to traditional algorithms.
-  • If the point cloud possesses an order, our method can facilitate the identification of the nearest point to $q$ among $\{p_{i}\}_{i=1}^{m}$, for any $m$, where $m<n$.
+  • If the point cloud possesses an order, our method can facilitate the identification of the nearest point to $q$ among $\{p_{i}\}_{i=1}^{k}$, for any $k$, where $k\leq n$.
 This approach can be employed to reduce the time complexity of Density Peak Clustering [9], thus enhancing its applicability in large-scale scenarios.
   • Based on our method, we developed a farthest point sampling algorithm with an average complexity of $O(k\log n)$ in 3D, where $n$ represents the total number of points and $k$ denotes the number of samples.
-Additionally, our algorithm exhibits significant potential.
-For instance, under various primitive types and distance metrics, as long as the incremental Voronoi or Delaunay construction algorithms for the primitives, as well as the distance query function from a point in space to the primitive, are provided, our nearest neighbor search framework can support nearest neighbor searches in these contexts.
-However, further development will be necessary in future work.
-II Releated Work
+II Related Work
 II-A Exact Nearest Neighbor Search
-For spatial data, the nearest neighbor problem is usually solved by building a spatial index [10].
-These indices can broadly be divided into two categories: those that partition the space and those that partition the data.
-Classic structures for space partitioning include KD Tree and quadtrees.
+For spatial data, the nearest neighbor problem is typically addressed by building a spatial indexing structure [10].
+These techniques can be broadly categorized into two types: those that partition space and those that partition data.
+Classic structures for space partitioning include KD Trees and quadtrees.
 A KD Tree [7] is a spatial partitioning structure that recursively divides the dataset along one of the $k$ coordinate system axes, typically alternating between them; for instance, with $k=2$, splits occur alternately on the x- and y-axes.
@@
 KD Trees can be extended to include objects beyond points, such as triangles or line segments.
-Determining the optimal splitting plane can be challenging in these cases; However, for points (or disks) in two dimensions, we can simply sort the points along each dimension and split at the median.
+Determining the optimal splitting plane can be challenging in these cases; however, for points (or disks) in two dimensions, one can simply sort the points along each dimension and split at the median.
 The construction of a KD Tree can be achieved in $O(n\log n)$ time, and it can find the nearest neighbors in $O(\log n)$ time.
-A octree is another spatial partitioning structure, but it specifically divides a 3D space into eight cubes or regions.
+An octree is another spatial partitioning structure that specifically divides 3D space into eight cubes.
 This is typically done recursively, with each subdivision occurring when the data within a region meets certain criteria, such as exceeding a maximum number of points or objects.
-Each internal node in the octree corresponds to a cube region of space, which is further subdivided into eight equal-sized cubes, represented by its children.
-The construction of an octree can be completed in $O(n\log n)$ time, and it can find the nearest neighbors in $O(\log n)$ time, too.
+Each internal node in the octree corresponds to a cube region of space, further subdivided into eight equal-sized cubes represented by its children.
+The construction of an octree can also be completed in $O(n\log n)$ time, and it can find the nearest neighbors in $O(\log n)$ time as well.
 The R Tree [8] is a classic structure for partitioning data.
-It is based on B-trees and shares a similar structure: both the number of objects in a leaf node and the number of children in an internal node are bounded by specified minimum and maximum limits.
-Additionally, all leaf nodes are at the same level.
-The R-tree can be built in $O(n\log n)$ time and allows for the identification of the nearest neighbors in $O(\log n)$ time.
-Grids and linear search methods are also classic approaches for nearest-neighbor searches.
-The grid-based approach improves efficiency by overlaying a regular grid on the entire simulation area and tracking the agents within each cell.
+It is based on B-trees and shares a similar structure: both the number of objects in a leaf node and the number of children in an internal node are bounded by specified minimum and maximum limits, and all leaf nodes are at the same level.
+The R Tree can be built in $O(n\log n)$ time and allows for the identification of nearest neighbors in $O(\log n)$ time.
+Grids and linear search methods are also traditional approaches for nearest-neighbor searches.
+The grid-based approach improves efficiency by overlaying a regular grid on the entire simulation area and tracking agents within each cell.
 During a query, the search includes the cell containing the query points and the adjacent cells.
 The construction time for this method is $O(n+m)$, where $m$ represents the number of cells in the grid.
-The worst-case query time is $O(n)$, but practical performance greatly depends on the distribution of agents and the size of the cells.
-On the other hand, linear search is suitable when the number of query points is very small.
+While the worst-case query time is $O(n)$, practical performance greatly depends on the distribution of agents and the size of the cells.
+Conversely, linear search is suitable when the number of query points is very small.
 This method finds the nearest points by calculating the distance from each target point to the query point.
-II-B Voronoi Diagram & Delaunay
-Suppose that we have a set of generators $\mathcal{S}=\{s_{i}\}_{i=1}^{n}$ in a given domain $\Omega$ that is equipped with a metric function $\mathcal{D}$, the Voronoi diagram involves partitioning $\Omega$ into regions such that the generator $s_{i}$ dominates a region
-$\text{Cell}(s_{i})=\{x\in\Omega~{}\big{|}~{}\mathcal{D}(s_{i},x)\leq\mathcal{D}(s_{j},x),j\neq i\}.$ (2)
-The definition of Voronoi diagrams may vary with domains, metrics, and generator types.
-The most prevalent version assumes that $\Omega$ represents Euclidean spaces, and the generators are exclusively points.
-As a fundamental tool, this has found widespread applications in computer graphics [3, 1, 11].
-The Delaunay is the dual structure of the Voronoi diagram, representing the adjacency relationships between the Voronoi cells.
+II-B Approximate Nearest Neighbor Search
+Graph-based algorithms are highly effective for approximate nearest neighbor (ANN) search due to their balance of accuracy and computational efficiency [11, 12, 13, 14].
+Typically, the process begins with constructing connections within a given point cloud, followed by iterative traversal of the graph starting from an entry point, which may be selected randomly or determined by a separate algorithm.
+At each step, the algorithm selects the closest neighboring node as the next base node, iteratively refining the current solution.
+The search terminates once a predefined stopping condition, such as a set number of distance evaluations, is met.
+One well-known graph structure used for search is the Navigable Small World (NSW, also known as Metricized Small World, MSW) [15, 14, 12].
+NSW constructs the graph by inserting elements in random order.
+Each element is then bidirectionally connected to its $M$ nearest neighbors among the previously inserted nodes.
+To improve search efficiency, many algorithms adopt a hierarchical strategy by creating multiple layers of the graph, which helps narrow the search space and accelerate the process.
+Our method adopts a similar graph-based strategy, emphasizing efficient connectivity and traversal.
+Unlike traditional graph-based approximate nearest neighbor search, our method precisely identifies the nearest points with strict guarantees of correctness.
+Furthermore, although we do not explicitly use hierarchical structures, our integrated construction and traversal strategies achieve comparable benefits, such as efficient narrowing of the search space and improved accuracy.
+II-C Voronoi Diagram & Delaunay
+Suppose we have a set of generators $\mathcal{S}=\{s_{i}\}_{i=1}^{n}$ in a given domain $\Omega$ equipped with a metric function $\mathcal{D}$.
+The Voronoi diagram partitions $\Omega$ into regions such that the generator $s_{i}$ dominates a region defined as
+$\text{Cell}(s_{i})=\{x\in\Omega~{}|~{}\mathcal{D}(s_{i},x)\leq\mathcal{D}(s_{j},x),\,\forall j\neq i\}.$
+The definition of Voronoi diagrams may vary based on domains, metrics, and types of generators.
+The most prevalent version assumes that $\Omega$ represents Euclidean spaces and that the generators are exclusively points.
+As a fundamental tool, Voronoi diagrams have found widespread applications in computer graphics [3, 1, 16].
+The Delaunay triangulation is the dual structure of the Voronoi diagram, representing the adjacency relationships between the Voronoi cells.
 Currently, Voronoi diagrams are typically computed by first calculating the Delaunay triangulation and then deriving the dual structure.
@@
 It is based on the empty circle property and works by incrementally inserting points and flipping triangles that do not satisfy this property.
-Common tools for computing Delaunay include TetGen [13] and CGAL [14], which can achieve an average computational complexity of $O(n\log n)$, with a worst-case complexity of $O(n^{2})$.
+Common tools for computing Delaunay triangulations include TetGen [18] and CGAL [19], which can achieve an average computational complexity of $O(n\log n)$ with a worst-case complexity of $O(n^{2})$.
 Voronoi diagrams can also be applied to nearest neighbor searches.
@@
 This property makes Voronoi diagrams highly relevant for solving nearest neighbor problems.
-However, the challenge lies in efficiently identifying the specific Voronoi cell containing $q$, as there currently isn’t an efficient algorithm for this task.
-This limitation restricts the practical use of Voronoi diagrams in nearest neighbor searches.
-II-C Farthest Point Sampling
+However, there is currently no efficient algorithm for identifying the specific Voronoi cell containing $q$.
+This restricts the practical use of Voronoi diagrams in nearest neighbor searches.
+II-D Farthest Point Sampling
 Farthest point sampling (FPS) is a technique widely employed in computational geometry and computer graphics applications.
@@
 FPS is particularly useful in applications requiring a representative subset of data points, such as mesh generation, surface simplification, or cluster initialization.
-However, a notable drawback of FPS is its computational cost.
-The complexity of the method is $O(kn)$, where $n$ is the total number of points in the dataset, and $k$ denotes the number of points to be sampled.
-This complexity can become prohibitive for large datasets or when a high number of samples is needed, making FPS less efficient in such cases.
+The time complexity of a typical FPS method [20, 21] is $O(kn)$, where $n$ is the total number of points in the dataset, and $k$ denotes the number of points to be sampled.
+The latest FPS method QuickFPS [22] organizes large-scale point clouds into multiple bins.
+By employing two mechanisms—merge computation and implicit computation on these bins—it significantly reduces external memory access and computation costs, thereby improving the speed of farthest point sampling.
 […]
@@
 III Methodology
-In this section, we introduce the key concept underlying the algorithms discussed in this paper, beginning with an exploration of linear search.
+In this section, we introduce the key concept underlying the algorithms, beginning with a solution of linear search.
 We assume that the given points are numbered from 1 to $n$.
-Given a query point $q$, we use $\Phi_{m}(q)$ to denote the closest point in $\{p_{i}\}_{i=1}^{m}$, where $m\leq n$.
+Given a query point $q$, we use $\Phi_{m}(q)$ to denote the closest point in the subset $\{p_{i}\}_{i=1}^{m}$, where $m\leq n$.
 This leads to the following state transfer equation:
-$\Phi_{m}(q)=\begin{cases}p_{1}&\text{if }m=1,\\ \Phi_{m-1}(q)&\text{if }m\neq 1\ \&\ \|q-\Phi_{m-1}(q)\|\leq\|q-p_{m}\|,\\ p_{m}&\text{if }m\neq 1\ \&\ \|q-\Phi_{m-1}(q)\|>\|q-p_{m}\|,\end{cases}$ (4)
+$\Phi_{m}(q)=\begin{cases}p_{1}&\text{if }m=1,\\ \Phi_{m-1}(q)&\text{if }m\neq 1\land\|q-\Phi_{m-1}(q)\|\leq\|q-p_{m}\|,\\ p_{m}&\text{if }m\neq 1\land\|q-\Phi_{m-1}(q)\|>\|q-p_{m}\|,\end{cases}$ (3)
 which can be understood as a dynamic programming problem, where $\Phi_{n}(q)$ is the actual nearest point to be extracted.
@@
 […]
-$\Phi_{m+k}(q)$ $=\Phi_{m+k-1}(q)=\cdots=\Phi_{m}(q).$
+$\mapsto\quad\Phi_{m+k}(q)$ $=\Phi_{m+k-1}(q)=\cdots=\Phi_{m}(q).$
 IV Algorithm
@@
 During the incremental construction of the Voronoi diagram, when $p_{j}$ is inserted, if the condition $p_{j}\vdash\text{Cell}(p_{i};\mathcal{V}_{j-1})$ is satisfied, we push $p_{j}$ to the end of $\mathcal{L}_{i}$.
-Query Table $\mathcal{L}$ is constructed synchronously as the Voronoi diagram is incrementally built.
+$\mathcal{L}$ is constructed simutaneously as the Voronoi diagram is incrementally built.
 Initially, we have only one site $p_{1}$, leaving its Query List empty.
@@
 The Query Table $\mathcal{L}$ is completed once all the sites are inserted.
-It is important to note that the order of elements stored in each $\mathcal{L}_{i}$ are consistent with the insertion order.
+It is important to note that the order of elements stored in each $\mathcal{L}_{i}$ is consistent with the insertion order.
 […]
@@
 The rightmost figure shows the query process.
-Since the Delaunay is the dual of the Voronoi diagram and reflects the adjacency relationships between Voronoi cells, to check whether $p_{m}\vdash\text{Cell}(p_{i};\mathcal{V}_{m-1})$, this can be further converted into checking whether $p_{m}$ and $p_{i}$ are adjacent in the Delaunay constructed from $\{p_{i}\}_{i=1}^{m}$, which is dual to $\mathcal{V}_{m}$.
+Since the Delaunay is the dual of the Voronoi diagram and encodes the adjacency relationships between Voronoi cells, to check whether $p_{m}\vdash\text{Cell}(p_{i};\mathcal{V}_{m-1})$, this can be further converted into checking whether $p_{m}$ and $p_{i}$ are adjacent in the Delaunay constructed from $\{p_{i}\}_{i=1}^{m}$, which is dual to $\mathcal{V}_{m}$.
 Algorithm 1 provides the pseudocode for the corresponding construction process.
@@
 This significantly reduces the computational cost of sampling.
-The remaining task is to dynamically track the Voronoi cells that contain the remaining points as the Voronoi diagram is incrementally constructed.
-This process is straightforward.
+The remaining task involves two parts: dynamically tracking the remaining points contained within the cells of the already inserted points during the incremental construction of the Voronoi diagram, and dynamically obtaining the next sampling point.
+The first part can be accomplished efficiently due to the local properties of Voronoi cells, while for the second part, we utilize a segment tree structure in our implementation.
 […]
@@
 Right: When a new point is inserted, only the distance of the uninserted points within the green and yellow cells may need to be updated.
-IV-D Complexity Analysis
+IV-D Incremental Delaunay Construction
+Currently, numerous incremental methods for Delaunay construction, such as parallel acceleration [24, 25] and pre-sorting points [26], are commonly used to speed up the process.
+However, due to the specific requirements of farthest point sorting and the need to generate a query table during the construction phase in our approach, traditional acceleration techniques for Delaunay construction are not directly applicable.
+During the incremental Delaunay construction process, one of the most time-consuming operations is locating the tetrahedron that contains a newly inserted point.
+According to Section IV-A and IV-B of our algorithm, when inserting point $p_{k}$, we have already constructed a query table for finding the nearest point within the set $\{p_{i}\}_{i=1}^{k-1}$.
+This query table stores information that allows for rapid identification of nearest points among the already inserted points.
+Leveraging this structure, we can quickly determine the nearest point to $p_{k}$.
+This nearest point then serves as the starting point for efficiently locating the tetrahedron containing $p_{k}$, significantly accelerating the Delaunay construction process.
+IV-E Complexity Analysis
 In this section, we will analyze the complexity of the farthest point sampling as a standalone component, rather than as part of the construction phase.
@@
 Therefore, the average space complexity of the Query Table is
-$O(\sum_{i=2}^{n}c)=c*O(n)=O(n),$ (7)
+$O(\sum_{i=2}^{n}c)=c\times O(n)=O(n),$ (6)
 where $n$ is the number of points.
@@
 Thus, the average time complexity of the nearest neighbor query phase is
-$O(\sum_{i=1}^{n-1}\frac{c}{i})=c*O(\log n)=O(\log n).$ (8)
+$O(\sum_{i=1}^{n-1}\frac{c}{i})=c\times O(\log n)=O(\log n).$ (7)
 Time Complexity of Farthest Point Sampling.
@@
 Thus, the additional complexity caused by the farthest point sampling process is:
-$O(\sum_{k=1}^{n-1}\left[\log(n-k)+\left(c*\frac{n-k}{k}\right)\right])$ (9)
-$=O(\sum_{k=1}^{n-1}\log(k)+c*n*\sum_{k=1}^{n-1}\frac{1}{k}-(n-1)*c)$ (10)
+$O(\sum_{k=1}^{n-1}\left[\log(n-k)+\left(c\times\frac{n-k}{k}\right)\right])$
+$=O(\sum_{k=1}^{n-1}\log(k)+cn\times\sum_{k=1}^{n-1}\frac{1}{k}-(n-1)\times c)$
 $=O(n\log n)$ (8)
@@
 In all the experiments below, Ours1 refers to our method using a random order for incremental Delaunay construction, while Ours2 indicates that the farthest point sorting strategy was applied.
-All the point cloud data used in the experiment has been included in the supplementary data.
 V-A Query Table
@@
 V-B Preprocessing Cost
-When a random sorting method is adopted, the preprocessing time of the current algorithm is equivalent to that of incremental Delaunay construction.
-However, at this stage, our preprocessing speed is slower than that of existing Delaunay construction libraries.
-This is primarily due to two factors.
-First, our method currently supports only serial execution.
-Second, certain algorithms, such as TetGen, reorder points before incremental Delaunay construction to enhance overall speed.
-However, this insertion order is not favorable for our queries.
-It is worth noting, however, that our method has the potential to accelerate preprocessing when integrated with Delaunay construction.
-In the current implementation, we present the preprocessing times of different algorithms across several classic models in Table II.
-Additionally, Figure 5 shows the preprocessing time with respect to the resolution of the Dragon data, comparing our method to other methods.
-It can be observed that at this stage, the preprocessing time of our method exceeds that of existing nearest neighbor search algorithms.
+The preprocessing time of our algorithm is primarily concentrated in the Delaunay construction phase, as it involves significant computational overhead in building and managing the complex graph structure.
+In this section, we present the preprocessing times of different algorithms across several classic models in Table II.
+Additionally, Figure 5 illustrates the preprocessing time relative to the resolution of the Dragon point cloud, providing a comparative analysis between our method and others.
+Furthermore, Figures 9 and Figure 10 show the preprocessing time of different methods on subsets of the ABC dataset [33] and the Thingi10k dataset [34].
+It can be observed that at this stage, our method currently exhibits longer preprocessing times compared to existing nearest neighbor search algorithms, particularly at higher resolutions, indicating areas that require further optimization.
 TABLE II: Preprocessing time ($s$) on classical models.
@@
 R-tree 0.005 0.014 0.022 0.062 0.170 0.243 0.311
-Ours1 1.097 3.205 4.692 14.46 41.58 62.13 72.31
-Ours2 1.876 5.818 8.473 35.64 96.11 263.3 307.6
+Ours1 0.188 0.703 0.984 3.531 10.359 15.219 18.985
+Ours2 0.265 0.765 1.406 4.782 15.157 22.953 26.938
 […]
-Figure 5: The preprocessing cost of KD Tree, R Tree and Ours on the Dragon data with varying resolutions.
-Top: preprocessing time.
-Bottom: the comparison about the preprocessing cost.
-Dashed line indicates where the ratio equals 1.
+Figure 5: The preprocessing cost of KD Tree, R Tree and Ours on the Dragon data with varying resolutions. Top: preprocessing time. Bottom: the comparison about the preprocessing cost.
+[…]
+Figure 6: Comparison about preprocessing cost on ABC dataset.
+Top: the time cost of preprocessing for KD Tree, R Tree and ours.
+Bottom: the comparison about the preprocessing cost among KD Tree, R Tree and ours.
+[…]
+Figure 7: Comparison about preprocessing cost on Thingi10K dataset.
+Top: the time cost of preprocessing for KD Tree, R Tree and ours.
+Bottom: the comparison about the preprocessing cost among KD Tree, R Tree and ours.
 V-C Query Performance
@@
 The query time are shown in Figures 9 and Figure 10.
-And the preprocessing time can be found in Figure 13 and Figure 14 in Appendix.
+And the preprocessing time can be found in Figure 6 and Figure 7.
 […]
@@
 Bottom: Dragon point cloud.
+V-G Farthest point sampling
+QuickFPS [22] is currently the fastest farthest point sampling strategy, achieving performance speedups of 230.5× over traditional CPU implementations, 43.4× over traditional GPU implementations, and 12.2× over previous state-of-the-art point cloud accelerators.
+It divides the point cloud into multiple buckets based on spatial location, and utilizes advanced techniques such as merged computation and implicit computation to accelerate the farthest point sampling process.
+Although the method proposed in this paper is not specifically designed for farthest point sampling, it also demonstrates impressive speed in this task.
+Using several traditional models as examples, with the farthest point sampling rate set at 25%, the sampling times of QuickFPS (CPU Version) and the method proposed in this paper are shown in Table IV.
+Moreover, compared to the complexity of QuickFPS, our proposed algorithm are both simpler and more intuitive.
+TABLE IV: Comparison of Sampling Times (s) between QuickFPS and Ours at a 25% Sampling Rate.
+Camel Bunny Dragon Kitten Armadillo Lucy Sponza
+Vertices 28934 72911 100313 291023 726367 1018219 1313504
+QuickFPS 0.016 0.094 0.156 1.109 2.391 2.875 3.969
+Ours 0.079 0.25 0.328 1.484 5.421 8.453 11.25
 VI Desirable Properties
-Beyond its speed advantages, we believe the current method holds significant potential, and future high-quality research could further expand the algorithm’s capabilities.
-In the following section, we will analyze several key aspects.
-VI-A Nearest Neighbor Among the First $n$ Points
+In light of its speed advantages, we believe that the current method harbors substantial potential for further advancements.
+In the following, we will delineate several pivotal aspects pertaining to its applications, aiming to elucidate the breadth and depth of its prospective utility in various domains.
+VI-A Nearest Neighbor Among the First $k$ Points
 One notable property of our algorithm is that, given any query point $q$ and integer $k$, we can efficiently identify the nearest point among $\{p_{i}\}_{i=1}^{k},k\leq n$.
@@
 Generally, computing $\{\delta_{i}\}_{i=1}^{n}$ requires calculating the distance between every pair of points, which has a complexity of $O(n^{2})$.
-However, for 3D points, using our algorithm, this problem can be solved in $O(n\log n)$ time.
+In contrast, our algorithm can efficiently solve this problem in $O(n\log n)$ time, assuming the data points are two-dimensional or three-dimensional.
+However, for higher-dimensional data, the time complexity of our algorithm increases due to the computationally expensive Delaunay triangulation process.
 VI-B Farthest Point Sampling.
-Farthest Point Sampling (FPS) is a widely used sampling algorithm due to its ability to ensure uniform sampling of points, making it broadly applicable.
-For instance, in the PointNet++ [25] framework for 3D point cloud deep learning, FPS is employed to sample and cluster points as part of the receptive field.
-The complexity of sampling $k$ points from $n$ points is $O(kn)$.
-Currently, there are approximate farthest point sampling algorithms designed for specific types of point clouds [26], but as far as we know, no breakthroughs have been made in reducing the complexity of exact farthest point sampling.
-However, based on our algorithm, we achieve a complexity of $O(k\log n)$ for sampling $k$ points from $n$ points in 3D space.
+Farthest Point Sampling (FPS) is a widely used sampling algorithm because it ensures uniform point distribution, making it highly applicable across various tasks.
+For instance, FPS is used in the PointNet++ [35] framework for 3D point cloud deep learning to ensure uniform sampling and clustering of points, which contributes to building an effective receptive field.
+Although the initial purpose of our method was not specifically for farthest point sampling, it naturally lends itself to a simple and fast FPS strategy.
+Compared to existing methods, our proposed FPS strategy is more straightforward and easier to implement, offering an intuitive solution for efficient point sampling.
 VI-C Various Primitives & Distance Metrics
@@
 Of course, we acknowledge that the incremental construction of Voronoi diagrams for various primitives is quite challenging; however, as outlined below, we also provide alternative options.
-VI-D Fast Preprocessing & Slow Query
-We acknowledge that the preprocessing efficiency of our algorithm is relatively low, especially in high-dimensional spaces.
-Furthermore, as previously mentioned, the incremental construction of Voronoi diagrams for various primitives under different distance metrics presents significant challenges.
-Therefore, we propose a new direction.
-As discussed in Section IV-A, when constructing the Query Table, we need to identify the set $\omega_{m}$ upon the insertion of $p_{m}$, as shown in Algorithm 1.
-The set is defined as follows:
+VI-D Fast Preprocessing with Compromising Query Performance
+In what follows, we aim to enhance the preprocessing efficiency at the expense of query performance.
+As elaborated in Section IV-A, during the construction of the Query Table, it is essential to identify the set $\omega_{m}$ upon inserting $p_{m}$, as illustrated in Algorithm 1.
+The set is defined by:
 $\omega_{m}=\{p_{i}~{}|~{}p_{m}\vdash\text{Cell}(p_{i};\mathcal{V}_{m-1})\}.$ (10)
-In fact, when dealing with complex high-dimensional or intricate scenarios, computing $\omega_{m}$ becomes extremely challenging.
-Therefore, we compute $\omega_{m}^{*}$ instead, which only needs to satisfy $\omega\subseteq\omega_{m}^{*}$, while still ensuring accurate nearest neighbor results.
-This approach offers a direction for performing fast nearest neighbor searches in high-dimensional spaces and scenarios involving different primitives and distance metrics.
-As a result, the computed set may contain redundancies, leading to a slight reduction in the speed of the nearest neighbor search.
-Additionally, we believe that even if $\omega_{m}$ does not belong to $\omega_{m}^{*}$, it could provide valuable insights for developing approximate nearest neighbor search algorithms.
-VI-E Fast Query & Slow Preprocessing
+Indeed, when confronted with complex high-dimensional or intricate scenarios, computing $\omega_{m}$ becomes exceedingly challenging.
+Consequently, we opt to compute $\omega_{m}^{*}$, which merely needs to satisfy $\omega_{m}\subseteq\omega_{m}^{*}$, while still ensuring accurate nearest neighbor results.
+This approach opens up avenues for executing rapid nearest neighbor searches in high-dimensional spaces and diverse primitive and distance metric contexts.
+As an inevitable outcome, the computed set may harbor redundancies, thereby causing a marginal decrease in the speed of the nearest neighbor search.
+Moreover, even if $\omega_{m}$ does not fall within $\omega_{m}^{*}$, it could furnish valuable insights conducive to the development of approximate nearest neighbor search algorithms.
+VI-E Fast Query with Elaborated Preprocessing
 Another encouraging finding is that the number of comparisons in our query process, as shown in Table I, can be further reduced, suggesting potential for additional algorithmic acceleration.
@@
 As shown in brown in (c).
-However, based on the query process, we can infer that $q\in\text{Cell}(p_{1};\mathcal{V}_{2})~{}\cap~{}\text{Cell}(p_{3};\mathcal{V}_{3})$.
-As shwon in (d).
-This enables us to design a more rigorous and precise occupancy evaluation.
+However, based on the query process, we can infer that $q\in\text{Cell}(p_{1};\mathcal{V}_{2})~{}\cap~{}\text{Cell}(p_{3};\mathcal{V}_{3})$, as is shown in (d).
+This enables us to design a more rigorous and precise evaluation criterion.
 Our method currently does not leverage this information during the query process, even though it can be obtained through complex preprocessing, which would, of course, involve substantial preprocessing overhead.
 […]
-Figure 12: (a) and (d): A simple incremental insertion process, with the red dot representing the query point $q$.
+Figure 14: (a) and (b): A simple incremental insertion process, with the red dot representing the query point $q$.
 (c) Based on the query process of (a)(b), we know the information $q\in\text{Cell}(p_{3};\mathcal{V}_{3})$, as shown in brown.
@@
 In brief, in addition to $q\in\text{Cell}(p_{i};\mathcal{V}_{i})$, we incorporated $q\in\text{Cell}(\Phi_{i-1};\mathcal{V}_{i-1})~{}\cap~{}\text{Cell}(p_{i};\mathcal{V}_{i})$.
-This region is used as the fundamental unit for occupancy evaluation, rather than a complete Voronoi cell.
+This region is used as the fundamental unit for occupancy evaluation, rather than a whole Voronoi cell.
 We were pleasantly surprised to find that, despite implementing only a simplified version, we achieved a significant reduction in the number of comparisons compared to Table I.
-As shown in Table IV, $Ours^{1}_{*}$ and $Ours^{2}_{*}$ represent the number of comparisons without and with farthest point sampling, respectively, under the strategy outlined in this section.
+As shown in Table V, $Ours^{1}_{*}$ and $Ours^{2}_{*}$ represent the number of comparisons without and with farthest point sampling, respectively, under the strategy outlined in this section.
 Remarkably, only around 100 pairwise distance comparisons are required to find the exact nearest neighbor for a query point among 1 million points.
@@
 Nevertheless, we believe this is a highly meaningful direction for further exploration.
-TABLE IV: The number of pairwise comparisons during the query process using different construction strategies.
+TABLE V: The number of pairwise comparisons during the query process using different construction strategies.
 Camel Bunny Dragon Kitten Armadillo Lucy Sponza
@@
 $\text{Ours}^{1}_{*}$ 93.218 97.49 108.545 107.708 127.760 134.123 124.615
-$\text{Ours}^{2}_{*}$ 75.453 74.055 88.089 82.812 94.953 106.269 124.547
+$\text{Ours}^{2}_{*}$ 75.453 74.055 88.089 82.812 94.953 106.269 108.203
 VII Limitations and Future Work
-The main drawback of our algorithm at present is its preprocessing cost, which limits its application in some scenarios.
-However, fortunately, our method supports nearest neighbor queries during the construction process, which may assist in the tetrahedron location problem during incremental Delaunay construction, offering a potential avenue for accelerating preprocessing.
-Therefore, in future work, we will focus on improving the preprocessing speed.
-Additionally, our method currently only supports nearest neighbor search; we aim to extend the current algorithm to support k-nearest neighbor queries in the future.
-Furthermore, all the properties mentioned in Section VI are goals we aim to address in our future work.
+The primary limitation of our algorithm is its high preprocessing cost, which restricts its applicability, particularly in large-scale scenarios or with high-dimensional datasets where efficiency is crucial.
+Therefore, in future work, we plan to couple the construction of our algorithm more closely with Delaunay computations, or alternatively, develop an approximate Delaunay construction to achieve faster preprocessing.
+These approaches aim to reduce the preprocessing burden and improve scalability for large-scale and high-dimensional applications.
+Additionally, our method currently only supports nearest neighbor search.
+In future work, we aim to extend the algorithm to support k-nearest neighbor queries, which would enhance its versatility.
+Moreover, addressing all the desirable properties outlined in Section VI remains a significant long-term goal for us.
 VIII Conclusion
-In this paper, we observe that when dealing with point clouds distributed over 2D manifolds in 3D space, or when query points are distant from the target points, the efficiency of traditional nearest neighbor search algorithms decreases.
+In this paper, we observe that traditional nearest neighbor search algorithms exhibit decreased efficiency when dealing with point clouds distributed over 2D manifolds in 3D space, or when query points are distant from the target points.
 In certain special cases, the complexity can degrade to $O(n)$.
-Inspired by the incremental Delaunay construction process, we propose a nearest neighbor search algorithm specifically for 2D maniflod point cloud.
-Our method achieves 1-10 times the speed of KD Tree in nearest neighbor searches.
+Inspired by the incremental Delaunay construction process, we propose a novel nearest neighbor search algorithm specifically designed for 2D manifold point cloud data.
+Our method achieves a speedup of 1-10 times compared to the KD Tree approach.
 We validated the efficiency advantages of our approach through extensive experiments.
+The high performance of our algorithm is also supported by evidence showing that it requires only about 100 pairwise distance comparisons to identify the point closest to the query point from a set of 1 million points.
 Additionally, our algorithm demonstrates significant potential.
@@
 Moreover, we have implemented a farthest point sampling algorithm with $O(k\log n)$ complexity in 3D space.
-Our mathod can efficiently identify the point closest to the query point from a set of 1 million points by performing about 100 pairwise distance comparisons.
 We believe that the algorithm presented in this paper holds significant potential for application and will make a substantial contribution to the community.
-[…]
-Figure 13: Comparison about preprocessing cost on ABC dataset.
-Top: the time cost of preprocessing for KD Tree, R Tree and ours.
-Bottom: the comparison about the preprocessing cost among KD Tree, R Tree and ours.
-[…]
-Figure 14: Comparison about preprocessing cost on Thingi10K dataset.
-Top: the time cost of preprocessing for KD Tree, R Tree and ours.
-Bottom: the comparison about the preprocessing cost among KD Tree, R Tree and ours.

### v5

@@
 More specifically, the DAG captures how the proximity structure evolves during the incremental construction of the Voronoi diagram of the data points.
-Experimental results demonstrate that our method achieves a 1x-10x speedup.
-Additionally, our algorithm offers several valuable features.
-For instance, it naturally supports an $O(k\log n)$ algorithm for farthest point sampling, where $k$ is the desired number of sample points.
-Moreover, density peak clustering, which involves finding the nearest point among the top K points, is typically considered to have a time complexity of $O(n^{2})$.
-With our algorithm, this can be reduced to $O(n\log n)$.
-We believe this work will inspire further research on the NNS problem.
+Experimental results demonstrate that our method achieves a 1-10x speedup.
+Additionally, our algorithm demonstrates significant practical value across diverse applications.
+We validated its effectiveness through extensive testing in four key applications: Point-to-Mesh Distance Queries, Iterative Closest Point (ICP) Registration, Density Peak Clustering, and Point-to-Segments Distance Queries.
+A particularly notable feature of our approach is its unique ability to efficiently identify the nearest neighbor among the first $k$ points in the point cloud—a capability that enables substantial acceleration in low-dimensional applications like Density Peak Clustering.
+As a natural extension of our incremental construction process, our method can also be readily adapted for farthest point sampling tasks.
+These experimental results across multiple domains underscore the broad applicability and practical importance of our approach.
 Index Terms:
-Nearest neighbor search, farthest point sampling, density peak clustering, delaunay, voronoi diragram.
+Nearest neighbor search, Delaunay triangulation, Voronoi diragram, farthest point sampling, density peak clustering,
 I Introduction
@@
 This approach can be employed to reduce the time complexity of Density Peak Clustering [9], thus enhancing its applicability in large-scale scenarios.
-  • Based on our method, we developed a farthest point sampling algorithm with an average complexity of $O(k\log n)$ in 3D, where $n$ represents the total number of points and $k$ denotes the number of samples.
+  • We demonstrate our algorithm’s practical effectiveness through comprehensive evaluation on four key applications: Point-to-Mesh Distance Queries, Iterative Closest Point Registration, Density Peak Clustering, and Point-to-Segments Distance Queries, showcasing its broad applicability in computational geometry scenarios.
 II Related Work
-II-A Exact Nearest Neighbor Search
+II-A Nearest Neighbor Search
 For spatial data, the nearest neighbor problem is typically addressed by building a spatial indexing structure [10].
@@
 Determining the optimal splitting plane can be challenging in these cases; however, for points (or disks) in two dimensions, one can simply sort the points along each dimension and split at the median.
-The construction of a KD Tree can be achieved in $O(n\log n)$ time, and it can find the nearest neighbors in $O(\log n)$ time.
+The construction of a KD Tree can be achieved in $O(n\log n)$ time, and it can find the nearest neighbors in $O(\log n)$ time on average [11].
 An octree is another spatial partitioning structure that specifically divides 3D space into eight cubes.
@@
 Each internal node in the octree corresponds to a cube region of space, further subdivided into eight equal-sized cubes represented by its children.
-The construction of an octree can also be completed in $O(n\log n)$ time, and it can find the nearest neighbors in $O(\log n)$ time as well.
+The computation cost for locating a leaf node in an octree is $O(depth)\approx O(\log n)$, where depth represents the average depth of the tree [12].
+However, nearest neighbor queries typically require additional traversal operations beyond simple leaf node access, potentially increasing the practical query time by either constant factors or additional complexity terms.
 The R Tree [8] is a classic structure for partitioning data.
 It is based on B-trees and shares a similar structure: both the number of objects in a leaf node and the number of children in an internal node are bounded by specified minimum and maximum limits, and all leaf nodes are at the same level.
-The R Tree can be built in $O(n\log n)$ time and allows for the identification of nearest neighbors in $O(\log n)$ time.
+The R*-tree [13] is an enhanced variant that incorporates a combined optimization of area, margin, and overlap of enclosing rectangles.
+It introduces forced reinsert to dynamically reorganize the structure and prevent unnecessary splits, demonstrating superior performance over existing R-tree variants across different data distributions and query types.
 Grids and linear search methods are also traditional approaches for nearest-neighbor searches.
 The grid-based approach improves efficiency by overlaying a regular grid on the entire simulation area and tracking agents within each cell.
-During a query, the search includes the cell containing the query points and the adjacent cells.
-The construction time for this method is $O(n+m)$, where $m$ represents the number of cells in the grid.
-While the worst-case query time is $O(n)$, practical performance greatly depends on the distribution of agents and the size of the cells.
-Conversely, linear search is suitable when the number of query points is very small.
-This method finds the nearest points by calculating the distance from each target point to the query point.
-II-B Approximate Nearest Neighbor Search
-Graph-based algorithms are highly effective for approximate nearest neighbor (ANN) search due to their balance of accuracy and computational efficiency [11, 12, 13, 14].
-Typically, the process begins with constructing connections within a given point cloud, followed by iterative traversal of the graph starting from an entry point, which may be selected randomly or determined by a separate algorithm.
-At each step, the algorithm selects the closest neighboring node as the next base node, iteratively refining the current solution.
-The search terminates once a predefined stopping condition, such as a set number of distance evaluations, is met.
-One well-known graph structure used for search is the Navigable Small World (NSW, also known as Metricized Small World, MSW) [15, 14, 12].
-NSW constructs the graph by inserting elements in random order.
-Each element is then bidirectionally connected to its $M$ nearest neighbors among the previously inserted nodes.
-To improve search efficiency, many algorithms adopt a hierarchical strategy by creating multiple layers of the graph, which helps narrow the search space and accelerate the process.
-Our method adopts a similar graph-based strategy, emphasizing efficient connectivity and traversal.
-Unlike traditional graph-based approximate nearest neighbor search, our method precisely identifies the nearest points with strict guarantees of correctness.
-Furthermore, although we do not explicitly use hierarchical structures, our integrated construction and traversal strategies achieve comparable benefits, such as efficient narrowing of the search space and improved accuracy.
-II-C Voronoi Diagram & Delaunay
-Suppose we have a set of generators $\mathcal{S}=\{s_{i}\}_{i=1}^{n}$ in a given domain $\Omega$ equipped with a metric function $\mathcal{D}$.
-The Voronoi diagram partitions $\Omega$ into regions such that the generator $s_{i}$ dominates a region defined as
-$\text{Cell}(s_{i})=\{x\in\Omega~{}|~{}\mathcal{D}(s_{i},x)\leq\mathcal{D}(s_{j},x),\,\forall j\neq i\}.$
-The definition of Voronoi diagrams may vary based on domains, metrics, and types of generators.
-The most prevalent version assumes that $\Omega$ represents Euclidean spaces and that the generators are exclusively points.
-As a fundamental tool, Voronoi diagrams have found widespread applications in computer graphics [3, 1, 16].
-The Delaunay triangulation is the dual structure of the Voronoi diagram, representing the adjacency relationships between the Voronoi cells.
+During query operations, the algorithm examines the cell containing the query point and its adjacent cells.
+However, this approach exhibits significant performance degradation in certain scenarios—particularly when data points are distributed on a spherical surface with the query point at the center.
+In such cases, the algorithm must traverse multiple cells in all directions, essentially performing a comprehensive dataset scan.
+Linear search, conversely, offers a straightforward solution by calculating distances between the query point and each target point, making it suitable only when dealing with very few query operations or small datasets.
+Additionally, sketching techniques have been proposed for similarity search acceleration.
+Mic et al.
+[14] introduced methods for selecting optimal sketches that transform objects into bit-strings.
+Their sketch construction often relies on reference points, or pivots, to partition the data space.
+While their approach focuses on Hamming embedding for approximate similarity filtering, our method constructs a ‘Query Table’ based on the incremental insertion of points, which begins the search from a fixed initial point.
+Our goal is the exact nearest neighbor search in metric spaces through Voronoi diagram construction.
+II-B Voronoi Diagram & Delaunay
+Given a set of generators $\mathcal{P}=\{p_{i}\}_{i=1}^{n}$ in a domain $\Omega$ equipped with a metric function $\mathcal{D}$, the Voronoi diagram partitions $\Omega$ into regions where each generator $p_{i}$ dominates a cell defined as:
+$\text{Cell}(p_{i})=\{x\in\Omega~|~\mathcal{D}(p_{i},x)\leq\mathcal{D}(p_{j},x),\,\forall j\neq i\}.$ (2)
+While Voronoi diagrams can be generalized across different domains, metrics, and generator types, the most commonly studied variant assumes Euclidean spaces with point generators.
+As fundamental geometric structures, Voronoi diagrams have found extensive applications in computer graphics [3, 1, 15].
+The Delaunay triangulation serves as the geometric dual of the Voronoi diagram, encoding the adjacency relationships between Voronoi cells.
+In 2D, the Delaunay triangulation consists of triangles whose vertices correspond to the Voronoi generators, while in 3D, it comprises tetrahedra with the same vertex correspondence.
+Figure 1 illustrates this duality in the 2D setting, where the Delaunay triangulation is depicted with black dashed lines and the Voronoi diagram with gray solid lines.
+This organizational structure makes it straightforward to obtain the adjacent sites of any given vertex, with many computational geometry libraries providing corresponding interfaces for such operations.
 Currently, Voronoi diagrams are typically computed by first calculating the Delaunay triangulation and then deriving the dual structure.
-The Watson algorithm [17] is a classic method for computing the Delaunay triangulation.
+The Bowyer–Watson algorithm [16] is a classic method for computing the Delaunay triangulation.
 It is based on the empty circle property and works by incrementally inserting points and flipping triangles that do not satisfy this property.
-Common tools for computing Delaunay triangulations include TetGen [18] and CGAL [19], which can achieve an average computational complexity of $O(n\log n)$ with a worst-case complexity of $O(n^{2})$.
-Voronoi diagrams can also be applied to nearest neighbor searches.
-By definition, if the closest point to a query point $q$ in $\mathbf{P}$ is $p_{k}$, then $q$ lies within the Voronoi cell of $p_{k}$.
-This property makes Voronoi diagrams highly relevant for solving nearest neighbor problems.
-However, there is currently no efficient algorithm for identifying the specific Voronoi cell containing $q$.
-This restricts the practical use of Voronoi diagrams in nearest neighbor searches.
-II-D Farthest Point Sampling
-Farthest point sampling (FPS) is a technique widely employed in computational geometry and computer graphics applications.
-It is designed to iteratively select points from a dataset such that each newly selected point is maximally distant from the previously chosen points.
-This method can be formally expressed as follows:
-$p_{i+1}=\arg\max_{p\in P}\left(\min_{p_{j}\in S}d(p,p_{j})\right),$ (2)
-where:
-  • $P$ represents the set of points in the point cloud,
-  • $S$ denotes the subset of points that have already been selected,
-  • $d(p,p_{j})$ is the distance between points $p$ and $p_{j}$,
-  • $p_{i+1}$ is the next point selected such that its minimum distance to any point in $S$ is maximized.
-This sampling strategy ensures that the selected points are well-distributed, making it effective for approximating shapes, surfaces, or distributions.
-FPS is particularly useful in applications requiring a representative subset of data points, such as mesh generation, surface simplification, or cluster initialization.
-The time complexity of a typical FPS method [20, 21] is $O(kn)$, where $n$ is the total number of points in the dataset, and $k$ denotes the number of points to be sampled.
-The latest FPS method QuickFPS [22] organizes large-scale point clouds into multiple bins.
-By employing two mechanisms—merge computation and implicit computation on these bins—it significantly reduces external memory access and computation costs, thereby improving the speed of farthest point sampling.
+Common tools for computing Delaunay triangulations include TetGen [17] and CGAL [18], which can achieve an average computational complexity of $O(n\log n)$ with a worst-case complexity of $O(n^{2})$ [19].
+[…]
+Figure 1: Example of Voronoi diagrams and Delaunay triangulations in 2D.
+The gray solid lines represent the Voronoi diagram partitioning the plane based on proximity to point generators (black dots), while the black dashed lines show the corresponding Delaunay triangulation that encodes the connectivity relationships between neighboring Voronoi cells.
+The geometric properties of Voronoi diagrams also enable nearest neighbor queries.
+By construction, if the closest point to a query location $q$ among generators $\mathbf{P}$ is $p_{k}$, then $q$ necessarily lies within $\text{Cell}(p_{k})$.
+However, the practical utility of this property is limited by the absence of efficient algorithms for point location within Voronoi cells, presenting an ongoing challenge for large-scale nearest neighbor applications.
 […]
@@
 (b) If the addition of the three sites (whose cells are colored in green) does not change the yellow cell, $\Phi_{m}(q)$ remains unchanged when $m$ is increased from 20 to 23.
-(b) When the 24-th site $p_{24}$ is added (see the cyan cell), the yellow cell diminishes; yet $\Phi_{24}(q)=\Phi_{23}(q)$ if $q$ is nearer to the old site.
+(c) When the 24-th site $p_{24}$ is added (see the cyan cell), the yellow cell diminishes; yet $\Phi_{24}(q)=\Phi_{23}(q)$ if $q$ is nearer to the old site.
 (d) $\Phi_{24}(q)=p_{24}$ if $q$ is nearer to the newly added site $p_{24}$.
@@
 This leads to the following state transfer equation:
-$\Phi_{m}(q)=\begin{cases}p_{1}&\text{if }m=1,\\ \Phi_{m-1}(q)&\text{if }m\neq 1\land\|q-\Phi_{m-1}(q)\|\leq\|q-p_{m}\|,\\ p_{m}&\text{if }m\neq 1\land\|q-\Phi_{m-1}(q)\|>\|q-p_{m}\|,\end{cases}$ (3)
+$\hskip-8.53581pt\Phi_{m}(q)=\begin{cases}p_{1}&\text{if }m=1,\\ \Phi_{m-1}(q)&\text{if }m\neq 1\land\|q-\Phi_{m-1}(q)\|\leq\|q-p_{m}\|,\\ p_{m}&\text{if }m\neq 1\land\|q-\Phi_{m-1}(q)\|>\|q-p_{m}\|,\end{cases}$ (3)
 which can be understood as a dynamic programming problem, where $\Phi_{n}(q)$ is the actual nearest point to be extracted.
@@
 […]
-Figure 2: We consider an incremental construction process of the Voronoi diagram of $\{p_{i}\}_{i=1}^{n}$.
-During the process, we append $p_{j}$ to $p_{i}$’s Query List if the addition of $p_{j}$ changes $p_{i}$’s cell.
-The rightmost figure shows the query process.
+Figure 3: We consider an incremental construction process for the Voronoi diagram and its corresponding Delaunay triangulation of $\{p_{i}\}_{i=1}^{n}$.
+During this process, we append $p_{j}$ to $p_{i}$’s Query List if the insertion of $p_{j}$ modifies $p_{i}$’s Voronoi cell.
+In other words, within the Delaunay triangulation, $p_{j}$ and $p_{i}$ are adjacent neighbors.
+The rightmost figure illustrates this query process.
 Since the Delaunay is the dual of the Voronoi diagram and encodes the adjacency relationships between Voronoi cells, to check whether $p_{m}\vdash\text{Cell}(p_{i};\mathcal{V}_{m-1})$, this can be further converted into checking whether $p_{m}$ and $p_{i}$ are adjacent in the Delaunay constructed from $\{p_{i}\}_{i=1}^{m}$, which is dual to $\mathcal{V}_{m}$.
+Figure 3 illustrates the synchronized incremental construction process of both Voronoi diagrams and their dual Delaunay triangulations.
+As Delaunay triangulations explicitly encode structural relationships between points, obtaining neighboring points becomes straightforward, with most computational geometry libraries providing efficient interfaces for this purpose.
 Algorithm 1 provides the pseudocode for the corresponding construction process.
@@
 Insert $p_{1}$, initialize the Delaunay structure $\mathbf{D}$.
-foreach $i$ in $[2,n]$ do
+[…]
   Initialize the set $\omega$ as empty.
@@
 Output: The closest point to $q$.
-getNearestNeighbor($p_{I}=p_{1}$)
+[…]
 foreach $p$ in $\mathcal{L}_{I}$ do
    if $\|q-p\|<\|q-p_{I}\|$ then
-           return getNearestNeighbor($p$)
+[…]
      end if
@@
 Algorithm 2 Nearest Neighbor Search
-IV-C Farthest Point Sampling
+[…]
+Figure 4: Coupling of Farthest Point sorting with incremental Voronoi Construction.
+(a): Uninserted points (gray) are distributed within the Voronoi cells of the inserted points (black).
+(b): When a new point is inserted, only the distance of the uninserted points within the green and yellow cells may need to be updated.
+IV-C Insertion Reordering Based on Farthest-Point Sorting
 Sections IV-A and  IV-B describe the algorithm for constructing the Query Table and performing nearest neighbor searches.
 The algorithm assumes a specific order of the point cloud and incrementally constructs the Voronoi / Delaunay based on this order.
-When the point cloud is unordered or there is no need to compute the nearest point to $q$ among the first $n$ points, we can sort the points in a specific manner to enhance the efficiency of the query phase.
-Intuitively, the more uniformly the inserted point cloud is distributed in space, the faster the query speed.
-This is because it facilitates a transition from large to small spatial jumps during the query process.
-Farthest point sampling, a commonly used downsampling method, aims to distribute the sampled points as uniformly as possible across the point cloud.
-Drawing inspiration from this, we set the number of sampled points equal to the total number of points in the cloud and use the sampling order to incrementally construct the Voronoi diagram.
+When the point cloud is unordered, we can sort the points in a specific manner to enhance the efficiency of either the preprocessing or query phase.
+Intuitively, the more uniformly the inserted point cloud is distributed in space, the faster the query speed, as it facilitates a transition from large to small spatial jumps during the query process.
+Farthest point sampling (FPS) is a technique that iteratively selects points from a dataset such that each newly selected point is maximally distant from the previously chosen points.
+This process can be formalized as:
+$p_{i+1}=\arg\max_{p\in P}\left(\min_{p_{j}\in S}d(p,p_{j})\right),$ (6)
+where $P$ represents the point cloud, $S$ denotes the already selected points, $d(p,p_{j})$ is the distance between points, and $p_{i+1}$ is the next selected point that maximizes its minimum distance to any point in $S$.
+This strategy ensures well-distributed point selection, making it effective for shape approximation and surface representation.
+Drawing inspiration from this, we set the number of sampled points equal to the total number of points in the cloud and use the sampling order to incrementally construct the Delaunay triangulation.
 Notably, by integrating this approach into our construction process, we developed a farthest point sampling strategy with reduced complexity, leading to faster sampling times.
@@
 The remaining task involves two parts: dynamically tracking the remaining points contained within the cells of the already inserted points during the incremental construction of the Voronoi diagram, and dynamically obtaining the next sampling point.
-The first part can be accomplished efficiently due to the local properties of Voronoi cells, while for the second part, we utilize a segment tree structure in our implementation.
+For the first task, we dynamically maintain a vector for each inserted point, storing the remaining points that are nearest to that point.
+We update these vectors dynamically during the incremental Delaunay construction process.
+For the second task, a segment tree [21] provides an excellent data structure that supports both single-point value updates within the interval and maximum value queries within same interval, both in logarithmic time.
+Therefore, during incremental Delaunay construction, we obtain all remaining points requiring distance updates through vector updates, update them in the segment tree (with inserted points set to negative infinity), and then leverage the segment tree’s logarithmic-time maximum query capability to retrieve the next farthest sampling point.
+IV-D Two Delaunay Construction Strategies
+For the construction of Delaunay triangulations, numerous algorithms exist; we employ the Bowyer–Watson algorithm [16], which performs incremental Delaunay construction through sequential point insertion.
+When inserting a new point, the algorithm first locates the Delaunay tetrahedron containing the point’s spatial location, then performs edge flipping operations based on the empty sphere property until the Delaunay criteria are satisfied.
+In our framework, different insertion orders for the Delaunay construction can achieve either faster construction speed or improved query performance.
+We introduce two strategies adopted in this work:
+Prioritizing Construction Speed.
+One time-consuming operation during Delaunay construction is locating the tetrahedron containing the newly inserted point.
+To accelerate Delaunay construction, a common approach involves pre-sorting the input point cloud using space-filling curves before performing incremental construction.
+This strategy ensures that the point insertion order exhibits spatial adjacency properties, enabling newly inserted points to quickly propagate from the previous insertion location to locate the tetrahedron containing the new point.
+Additionally, this approach provides excellent spatial locality during the construction process.
+We adopt the point cloud sorting strategy from CGAL [22], specifically the spatial_sort function, which first organizes the point set into random buckets of increasing sizes, with Hilbert sorting applied only within each bucket.
+Our experiments demonstrate that this strategy achieves fast Delaunay construction speed with minimal impact on query performance.
+Prioritizing Query Performance.
+Farthest Point Sorting ensures that inserted points are distributed as uniformly as possible throughout the full point cloud, which intuitively leads to faster query speeds—a hypothesis validated by our experiments.
+The farthest point sampling strategy has been detailed in Section IV-C.
+Here, we primarily focus on how our algorithm can be integrated to accelerate the Delaunay construction process under the farthest point sampling ordering.
+Our algorithm possesses an important property: as illustrated in the Figure 3 (b,c), before inserting the 5th point, the query table for the first four points has already been constructed (shown in the bottom portion of the Figure 3 (b) ), enabling queries for the nearest point among the first 4 points.
+Consequently, when inserting point $p_{k}$, we can already utilize Algorithm 2 to query the nearest point to $p_{k}$ within $\{p_{i}\}_{i=1}^{k-1}$.
+As previously mentioned, one time-consuming operation during Delaunay construction is locating the tetrahedron containing the newly inserted point.
+Therefore, when inserting the $k$-th point, we can quickly determine the nearest point to $p_{k}$ from $\{p_{i}\}_{i=1}^{k-1}$.
+This nearest point then serves as the starting location for efficiently locating the tetrahedron containing $p_{k}$, significantly accelerating the Delaunay construction process.
+IV-E Time Complexity Analysis
+Preprocessing.
+Our preprocessing phase involves the incremental construction of the Delaunay triangulation, during which we store the requisite information for subsequent queries.
+Consequently, the preprocessing time complexity aligns with that of Delaunay triangulation construction.
+On average, the time complexity for constructing a Delaunay triangulation in 3D is $O(n\log n)$ [23], while the worst-case time complexity is $O(n^{2})$ [19].
+However, this worst-case scenario is rarely encountered in typical construction processes.
+Querying.
+Due to the difficulty in evaluating the impact of point cloud distribution on Delaunay construction, providing a rigorous theoretical proof for average-case complexity remains challenging.
+Therefore, we present an experimental average complexity based on empirical observations, followed by a theoretical worst-case complexity analysis.
+The experimental section validates the correctness of our experimental average complexity analysis.
+Experimental Average Complexity
+When querying the nearest neighbor for point $q$, the complexity of the query phase depends on the number of times the Voronoi cell containing $q$ undergoes shrinkage during the incremental construction process.
+Each time the Voronoi cell containing $q$ shrinks due to the insertion of a new point, a comparison operation is required to potentially update the nearest neighbor.
+More specifically, assume $\Phi_{m}(q)$ represents the current nearest neighbor of $q$ among the first $m$ points, and a new point $p_{m+1}$ is inserted.
+Based on empirical observations, on average, the Voronoi cell of $p_{m+1}$ is adjacent to a constant number of existing cells, denoted as $c$ [24].
+Consequently, the probability that $p_{m+1}$ is adjacent to $\Phi_{m}(q)$ in the Delaunay triangulation is approximately $c/m$, which corresponds to the likelihood of requiring a comparison operation.
+Thus, the average time complexity of the nearest neighbor query phase is:
+$O\left(\sum_{i=1}^{n-1}\frac{c}{i}\right)=c\times O(\log n)=O(\log n).$ (7)
+Theoretical Worst-Case Complexity
+The theoretical worst-case complexity of our algorithm is $O(n)$, as it relies on pairwise point comparisons and eliminates at least one invalid candidate with each comparison.
+A typical example is when all points in 3D space lie on a straight line, ordered sequentially as $p_{1},p_{2},\ldots,p_{n}$, with Delaunay construction following this insertion order.
+In this degenerate configuration, each point $p_{k}$ encroaches only upon $p_{k-1}$, creating a chain-like structure that forces linear traversal during nearest neighbor queries, thereby degrading the algorithm to $O(n)$ complexity.
 […]
-Figure 3: Coupling of Farthest Point sorting with incremental Voronoi Construction.
-Left: Uninserted points (gray) are distributed within the Voronoi cells of the inserted points (black).
-Right: When a new point is inserted, only the distance of the uninserted points within the green and yellow cells may need to be updated.
-IV-D Incremental Delaunay Construction
-Currently, numerous incremental methods for Delaunay construction, such as parallel acceleration [24, 25] and pre-sorting points [26], are commonly used to speed up the process.
-However, due to the specific requirements of farthest point sorting and the need to generate a query table during the construction phase in our approach, traditional acceleration techniques for Delaunay construction are not directly applicable.
-During the incremental Delaunay construction process, one of the most time-consuming operations is locating the tetrahedron that contains a newly inserted point.
-According to Section IV-A and IV-B of our algorithm, when inserting point $p_{k}$, we have already constructed a query table for finding the nearest point within the set $\{p_{i}\}_{i=1}^{k-1}$.
-This query table stores information that allows for rapid identification of nearest points among the already inserted points.
-Leveraging this structure, we can quickly determine the nearest point to $p_{k}$.
-This nearest point then serves as the starting point for efficiently locating the tetrahedron containing $p_{k}$, significantly accelerating the Delaunay construction process.
-IV-E Complexity Analysis
-In this section, we will analyze the complexity of the farthest point sampling as a standalone component, rather than as part of the construction phase.
-Time Complexity of Query Table Construction.
-The complexity of the Query Table Construction is the same as that of the incremental Delaunay construction.
-On average, the time complexity for constructing a Delaunay in 3D is $O(n\log n)$, while the worst-case time complexity is $O(n^{2})$.
-However, this worst-case scenario is seldom encountered in typical construction processes.
-Space Complexity of the Query Table.
-On average, the number of neighbors for a Voronoi cell in 3D is a constant, denoted as $c$.
-This implies that when a new point is inserted, it is adjacent to $c$ cells on average, requiring the point to be appended to the end of $c$ Query Lists.
-Therefore, the average space complexity of the Query Table is
-$O(\sum_{i=2}^{n}c)=c\times O(n)=O(n),$ (6)
-where $n$ is the number of points.
-Time Complexity of Nearest Neighbor Query.
-When querying the nearest neighbor for the point $q$, the complexity of the query phase depends on the number of times the cell containing $q$ (which may change during the query process) shrinks during the incremental construction, as a comparison is needed each time the Voronoi cell containing $q$ shrinks.
-Simply put, assuming $\Phi_{m}(q)$ is known and a new point $p_{m+1}$ is inserted, the cell of $p_{m+1}$ typically adjacent to $c$ cells on average.
-Therefore, the probability that $\text{Cell}(\Phi_{m}(q);\mathcal{V}_{m})$ shrinks is approximate $c/m$, indicating the possibility of needing a comparison.
-Thus, the average time complexity of the nearest neighbor query phase is
-$O(\sum_{i=1}^{n-1}\frac{c}{i})=c\times O(\log n)=O(\log n).$ (7)
-Time Complexity of Farthest Point Sampling.
-When we have already inserted $k$ points to construct the Voronoi diagram, each cell, on average, contains ${(n-k)}/{k}$ remaining points.
-Given the distances from all remaining points to the nearest sampled points are known, selecting the next point for sampling has a complexity of $O(\log(n-k))$.
-Knowing that the newly inserted point is adjacent to $c$ cells, the number of remaining points that potentially need their distances updated is $c\times(n-k)/k$.
-Thus, the additional complexity caused by the farthest point sampling process is:
-$O(\sum_{k=1}^{n-1}\left[\log(n-k)+\left(c\times\frac{n-k}{k}\right)\right])$
-$=O(\sum_{k=1}^{n-1}\log(k)+cn\times\sum_{k=1}^{n-1}\frac{1}{k}-(n-1)\times c)$
-$=O(n\log n)$ (8)
+Figure 5: Representative 3D models from the ABC dataset [25].
+[…]
+Figure 6: Representative 3D models from the Thingi10K dataset [26].
+[…]
+Figure 7: Representative 3D models from the MedShapeNet dataset [27].
 V Results
-Hardware environment. Our algorithm was implemented in C++ on a platform featuring a 3.4 GHz AMD Ryzen 9 5950X 16-Core CPU, 64GB of RAM, and running the Windows 11 operating system.
-Experimental parameters.
-We utilized CGAL [19] for Delaunay computations.
-In our experiments, we used random sorting and farthest point sorting as insertion order.
-We measured and recorded both the construction performance and query performance for each method.
-The space for randomly generating query points was confined within the bounding box with 8x the volume (with edge lengths twice that of the original bounding box) of the input point cloud.
-The number of query points generated was set at 1000K.
+Platform.
+Our algorithm was implemented in C++ and evaluated on a Mac mini equipped with an M4 CPU, 16GB of RAM, running the macOS operating system.
+We utilized CGAL [18] for Delaunay triangulation computations and employed CGAL’s multi-scale sorting approach and farthest point sorting as the insertion order strategies.
 Relevant approaches for comparison.
-Currently, there are numerous algorithms [27, 28, 29, 12] available for nearest neighbor search.
-In this paper, we selected two of the fastest and most classic methods for comparison.
+We selected six classic methods as comparison baselines, and together with our proposed method (incorporating two different pre-sorting strategies), we evaluated a total of eight approaches.
   1. 1.
-    KD Tree: We utilized the implementation from the nanoflann library [30], setting the maximum number of leaf nodes to 10.
-According to tests referenced in [31], this implementation is currently the fastest for nearest neighbor searches.
+    KD Tree: From the nanoflann library [28], with the maximum number of leaf nodes set to the default value of 10, which achieves nearly optimal query performance according to the benchmark results provided by the official repository.
   2. 2.
-    R Tree: We used the implementation from Boost Library [32].
-In all the experiments below, Ours1 refers to our method using a random order for incremental Delaunay construction, while Ours2 indicates that the farthest point sorting strategy was applied.
+    R* Tree: From Boost library [29], with the maximum number of elements per node set to 10, consistent with the KD Tree configuration.
+  3. 3.
+    Octree: From PCL library [30], with the resolution parameter set to 3 times the inter-point distance.
+  4. 4.
+    BD Tree: From ANN library, with the maximum number of elements in leaf node set to 10, and we only use it for exact nearest neighbor queries.
+  5. 5.
+    Voronoi: From CGAL library, utilizing its provided nearest_vertex function for nearest neighbor queries.
+  6. 6.
+    Grid: Our own implementation, with cell side length set to 3 times the inter-point distance.
+  7. 7.
+    Ours1: Our method, using CGAL’s multi-scale sorting method spatial_sort for incremental Delaunay construction.
+  8. 8.
+    Ours2: Our method, using farthest point sorting strategy for incremental Delaunay construction.
+Additionally, it is worth noting that the Grid method exhibits slower performance when there is a considerable distance between query points and target points.
+The Voronoi-based method is also relatively slow.
+Therefore, we only include results from these two methods in certain experiments.
+Datasets and Query Point Generation.
+We selected three datasets and several standard models for algorithm evaluation.
+For nearest neighbor queries, we utilize the vertices of these 3D models as the point cloud data.
+After filtering out models that are excessively large or small, the datasets cover the following ranges:
+  1. 1.
+    ABC Dataset [25]: We tested on a subset of 10,000 models, with the final dataset covering 600 to 600,000 points. Figure 5 shows representative models from this dataset.
+  2. 2.
+    Thingi10K Dataset [26]: Originally containing 10,000 models, the filtered dataset covers 400 to 140,000 points. Figure 6 displays sample models from this collection.
+  3. 3.
+    MedShapeNet Dataset [27]: We evaluated on a subset of 10,000 models, with the processed dataset covering 200 to 150,000 points. Figure 7 illustrates representative models from this dataset.
+For query point generation, we randomly distribute 1 million query points within the 2× axis-aligned bounding box for each test case and perform nearest neighbor queries.
+[…]
+Figure 8: The length of the Query List for different index points. Top: CGAL’s multi-scale sorting. Bottom: farthest point sorting.
+[…]
+Figure 9: Query process convergence under two insertion strategies. Average nearest distance (normalized by point cloud diagonal) versus number of comparisons, averaged over one million queries.
+TABLE I: The average (Avg) and variance (Var) of the length of the query list and the number of comparisons required per query on classical models.
+Camel Bunny Dragon Kitten Armadillo Lucy Sponza
+Vertices 28934 72911 100313 291023 726367 1018219 1313504
+$\text{Ours}^{1}$ Avg 16.753 16.445 16.371 15.508 15.864 16.106 15.765
+Var 20.099 19.281 16.619 19.181 19.026 19.301 18.713
+Tested 188.0 183.96 240.15 214.39 274.27 304.22 339.05
+$\text{Ours}^{2}$ Avg 14.414 14.340 14.197 14.115 14.246 14.333 13.638
+Var 16.783 15.794 16.355 16.207 15.947 16.016 15.927
+Tested 144.87 136.28 162.68 152.35 193.15 211.39 235.85
+[…]
+Figure 10: Preprocessing time comparison across ABC, Thingi10K, and MedShapeNet datasets (from left to right).
+[…]
+Figure 11: Preprocessing cost comparison of various methods on the Dragon model across different resolutions.
+Top: absolute preprocessing time.
+Middle and Bottom: relative preprocessing cost ratios between methods.
 V-A Query Table
-Length of the query list for a vertex
-Recall that the Query Table maintains a list of points encroached upon during the incremental construction of the Delaunay for each vertex in $V$.
-It is evident that the Voronoi cells of points inserted earlier often shrink more times, resulting in longer Query Lists.
-Therefore, to investigate how the length of the Query List varies with different insertion indexes, we conducted a study using the $100$K-vertex dragon point cloud as an example.
+Analysis of Query List Length.
+Recall that each vertex maintains a Query List, which is a vector storing the subsequently inserted points that become adjacent to the current vertex in the Delaunay triangulation during the incremental construction process.
+It is evident that points inserted earlier experience more subsequent insertions, resulting in longer Query Lists as they accumulate more neighbors over time.
+Therefore, to investigate how the length of the Query List varies with different insertion indexes, we conducted a study using the 100100100K-vertex dragon point cloud as an example.
 As shown in Figure 8, we have plotted the length of the query list for each point with respect to its insertion order.
-Although the impact of random sorting and farthest point sorting strategies on the length of the Query List is not readily apparent from the figure, we calculated the average lengths of the Query Lists to be 15.684 and 14.165, with standard deviations of 17.705 and 16.491, respectively.
-The farthest point sorting resulted in a shorter and more uniform distribution of the Query List.
-Similar results were obtained across multiple data.
-This characteristic enables the query table generated through farthest point sorting to facilitate faster nearest-neighbor searches.
-[…]
-Figure 4: The length of the Query List for different index points. Top: random sorting. Bottom: farthest point sorting.
-Number of tested vertices
+The results reveal that two different insertion orders lead to distinctly different query length distribution trends.
+CGAL’s multi-scale sorting method exhibits a clear two-stage characteristic: the first stage demonstrates significantly longer average query lengths compared to the second stage, while the query lengths fluctuate around their respective average values within each stage.
+In contrast, the query lengths obtained through farthest point sampling show a steady declining trend.
+Number of Comparisons.
 As shown in Algorithm 2, the nearest neighbor search problem is transformed into the task of calculating distances between points.
@@
 It can be observed that, after utilizing the farthest point sorting method, both the average length (Avg) and variance (Var) of the Query Lists decreased, accompanied by a reduction in the average number of comparisons during the nearest neighbor search process.
-TABLE I: The average (Avg) and variance (Var) of the length of the query list and the number of comparisons required per query on classical models.
+Query Process Convergence.
+In accordance with Algorithm 2, our method iteratively evaluates target points during query execution while dynamically updating the current nearest distance estimate.
+To quantitatively assess the influence of varying insertion orders on query efficiency, we conducted a comprehensive evaluation using the 100K dragon model.
+Specifically, we generated 1 million query points within the search space and executed nearest neighbor queries for each point.
+During execution, we systematically recorded the current nearest distance at each comparison step, subsequently computing the average distance across all queries per step.
+Figure 9 demonstrates the convergence behavior of this average nearest distance (normalized as a percentage of the point cloud’s diagonal length) as a function of comparison count under two distinct insertion ordering strategies.
+The results reveal that the Farthest Point Sampling strategy facilitates substantially accelerated and more consistent convergence throughout the query process, empirically validating its efficacy in enhancing search performance.
+TABLE II: Preprocessing time ($s$) on classical models
 Camel Bunny Dragon Kitten Armadillo Lucy Sponza
 Vertices 28934 72911 100313 291023 726367 1018219 1313504
-$\text{Ours}^{1}$ Avg 15.913 15.805 15.684 15.397 15.832 16.034 14.631
-Var 17.911 17.264 17.705 18.834 17.631 17.968 20.431
-Tested 171.783 166.726 209.516 182.719 234.742 276.549 418.285
-$\text{Ours}^{2}$ Avg 14.414 14.345 14.165 13.964 14.370 14.478 13.067
-Var 16.783 15.906 16.491 17.126 16.172 16.584 19.681
-Tested 144.754 135.831 161.822 155.806 197.042 235.085 358.413
+KD-tree 0.003 0.006 0.009 0.025 0.068 0.098 0.123
+R-tree 0.004 0.007 0.010 0.028 0.078 0.109 0.170
+Octree 0.002 0.003 0.004 0.011 0.033 0.047 0.055
+BD-tree 0.005 0.010 0.014 0.044 0.118 0.163 0.198
+Grid 0.002 0.004 0.008 0.024 0.313 0.362 0.053
+Voronoi 0.057 0.143 0.189 0.517 1.359 1.943 2.466
+Ours1 0.072 0.182 0.249 0.685 1.736 2.481 3.170
+Ours2 0.103 0.327 0.466 1.698 5.005 7.454 9.602
 V-B Preprocessing Cost
@@
 In this section, we present the preprocessing times of different algorithms across several classic models in Table II.
-Additionally, Figure 5 illustrates the preprocessing time relative to the resolution of the Dragon point cloud, providing a comparative analysis between our method and others.
-Furthermore, Figures 9 and Figure 10 show the preprocessing time of different methods on subsets of the ABC dataset [33] and the Thingi10k dataset [34].
+Figure 11 further illustrates how preprocessing time scales with point cloud resolution using the Dragon model as a representative example.
+We extend this evaluation to larger datasets in Figure 10, which demonstrates preprocessing performance across the ABC [25], Thingi10k [26], and MedShapeNet [27] datasets.
 It can be observed that at this stage, our method currently exhibits longer preprocessing times compared to existing nearest neighbor search algorithms, particularly at higher resolutions, indicating areas that require further optimization.
-TABLE II: Preprocessing time ($s$) on classical models.
+V-C Query Performance
+Query efficiency constitutes the primary advantage of our proposed approach.
+To systematically evaluate this performance benefit, we conducted a multi-faceted analysis across various datasets and resolutions.
+Using the Dragon dataset as a controlled benchmark, we measured query costs with respect to point cloud resolution, as illustrated in Figure 12.
+This analysis demonstrates a consistent and significant performance advantage for our method across all tested resolutions.
+Table III complements this analysis by providing detailed query performance comparisons across a diverse set of classical models at their standard resolutions, further validating our method’s superior efficiency.
+To establish the generalizability of these performance advantages, we expanded our evaluation to large-scale datasets with diverse geometric characteristics: Thingi10K [26], ABC [25], and MedShapeNet [27].
+Figure 13 summarizes these comprehensive results, revealing consistent performance patterns across all tested repositories.
+Notably, our method achieves query acceleration factors of 1-10× compared to current state-of-the-art approaches, with substantial improvements observed across various model types.
+V-D Memory Usage
+Space complexity represents another critical dimension for algorithm evaluation.
+To provide a comprehensive assessment, we analyzed the memory requirements of our approach alongside existing methods.
+The primary memory consumption in our algorithm stems from three components: incremental Delaunay triangulation construction, Query Table storage, and segment tree structures when implementing farthest point sampling.
+It should be noted that precise memory quantification presents inherent challenges due to system-level memory management mechanisms, resulting in measurement fluctuations.
+For statistical robustness, our analysis excludes data points with memory utilization below 0.2M.
+Figure 14 presents the memory consumption of our algorithm across the ABC, Thingi10K, and MedShapeNet datasets.
+The empirical results demonstrate that while our method exhibits moderately increased memory requirements compared to traditional approaches, this additional memory overhead remains within acceptable bounds given the significant query performance advantages.
+This trade-off between space complexity and query efficiency represents a favorable compromise for applications where query speed is paramount.
+[…]
+Figure 12: Average query time on the Dragon model across varying resolutions.
+Top: Absolute query time ($\mu$s) for different methods.
+Middle and Bottom: Performance ratios between methods, with dashed line indicating equal performance (ratio = 1).
+[…]
+Figure 13: Query time comparison across ABC, Thingi10K, and MedShapeNet datasets (from left to right). Dashed line indicates equal performance (ratio = 1).
+[…]
+Figure 14: Memory utilization comparison across ABC, Thingi10K, and MedShapeNet datasets.
+TABLE III: Average query time ($\mu$s) comparison across classical models. The highest and second-fastest query times are highlighted in bold and underlined, respectively
 Camel Bunny Dragon Kitten Armadillo Lucy Sponza
 Vertices 28934 72911 100313 291023 726367 1018219 1313504
-KD-tree 0.004 0.011 0.016 0.051 0.149 0.215 0.281
-R-tree 0.005 0.014 0.022 0.062 0.170 0.243 0.311
-Ours1 0.188 0.703 0.984 3.531 10.359 15.219 18.985
-Ours2 0.265 0.765 1.406 4.782 15.157 22.953 26.938
+KD-tree 1.109 2.542 1.585 5.402 6.003 4.296 1.815
+R*-tree 1.450 2.584 1.808 4.974 4.967 3.616 1.446
+Octree 5.060 8.475 5.719 17.06 17.21 14.89 10.22
+BD-tree 2.344 4.721 5.715 13.11 78.04 23.02 2.674
 […]
-Figure 5: The preprocessing cost of KD Tree, R Tree and Ours on the Dragon data with varying resolutions. Top: preprocessing time. Bottom: the comparison about the preprocessing cost.
+Voronoi 6.815 9.965 5.692 27.01 26.48 39.10 36.92
 […]
-Figure 6: Comparison about preprocessing cost on ABC dataset.
-Top: the time cost of preprocessing for KD Tree, R Tree and ours.
-Bottom: the comparison about the preprocessing cost among KD Tree, R Tree and ours.
+Ours2 0.201 0.235 0.213 0.322 0.377 0.401 0.617
+V-E Performance on Uniform Distributions
+To evaluate our algorithm’s efficiency across different data distributions, we tested performance on uniform point clouds.
+Unlike previous experiments, we generated query points within exactly the same spatial bounds as the target points—not in an expanded bounding box.
+As shown in Figure 15, which presents both query and preprocessing times, our method achieves query efficiency comparable to optimized KD Tree and Grid implementations while outperforming other approaches.
+These results demonstrate our algorithm’s effectiveness across diverse point distribution patterns beyond its primary optimization for manifold data.
 […]
-Figure 7: Comparison about preprocessing cost on Thingi10K dataset.
-Top: the time cost of preprocessing for KD Tree, R Tree and ours.
-Bottom: the comparison about the preprocessing cost among KD Tree, R Tree and ours.
-V-C Query Performance
-Using the Dragon dataset as an example, we present the query cost of our method compared to other approaches with respect to resolution in Figure 8, demonstrating the significant advantage of our method in this scenario.
-In Table III, we provide more detailed examples for several classical models.
+Figure 15: Performance evaluation on uniform point clouds with varying numbers of target points.
+Top: Query time comparison shows our method’s efficiency comparable to KD Tree and Grid while outperforming others.
+Bottom: Preprocessing time comparison.
+All query points were generated within the exact same spatial bounds as the target points.
+V-F Near and Far Query Points.
+The spatial relationship between query points and target points significantly influences nearest neighbor search efficiency.
+To quantify this effect, we conducted experiments varying the query space extent relative to the target point cloud.
+We tested three distinct distributions: uniform point clouds, Earth city models, and the Dragon dataset, each containing 1 million points.
+Figure 16 presents query performance as a function of spatial scaling.
+The horizontal axis represents the ratio between query generation space and target point bounding box (e.g., a value of 2 indicates query points generated within a 2× axis-aligned bounding box).
+Corresponding preprocessing times are provided in Table IV.
+The results reveal an interesting pattern: even in uniform distributions, where our method shows comparable performance when query points share the same bounds as target points, our approach gains significant advantages as soon as the query space extends slightly beyond the target bounds.
+For non-uniform distributions like Earth city models and Dragon dataset, our method consistently outperforms traditional approaches across all spatial configurations.
 […]
-Figure 8: Average query time on the Dragon model with varying resolutions.
-Top: Average query time ($\mu$s) among KD tree, R tree and ours.
-Bottom: The comparison about the average time different methods.
-Dashed line indicates where the ratio equals 1.
-TABLE III: Average query time ($\mu s$) of KD Tree, R Tree, and Ours on classical models. The highest and second-fastest query times are highlighted in bold and underlined, respectively.
-Camel Bunny Dragon Kitten Armadillo Lucy Sponza
-Vertices 28934 72911 100313 291023 726367 1018219 1313504
-KD Tree 1.639 3.795 2.475 8.475 12.81 9.796 4.407
-R Tree 1.623 2.739 2.034 5.359 5.241 4.518 1.851
-Ours1 0.423 0.559 0.540 0.957 1.132 1.293 2.146
-Ours2 0.344 0.482 0.388 0.752 0.834 1.070 1.694
-V-D Tests on ABC & Thingi10K
-Thingi10K [34] and ABC [33] are two large-scale 3D datasets that contain diverse models.
-To conduct a comprehensive comparison among KD Tree, R Tree, and Ours, we ran them on the dataset to compare both the preprocessing cost and the query cost.
-The query time are shown in Figures 9 and Figure 10.
-And the preprocessing time can be found in Figure 6 and Figure 7.
+Figure 16: Average query time with varying sampling box sizes.
+The horizontal axis shows the scaling factor applied to each dimension of the target point bounding box to define the query generation space.
+Results for uniform point cloud (top), Dragon model (middle), and Earth city data (bottom).
+TABLE IV: Preprocessing time comparison for uniform point cloud, Dragon model, and Earth city data (1M points each).
+Octree BD Tree R* Tree KD Tree Ours1 Ours2
+Uniform 0.047 0.145 0.121 0.105 2.317 6.775
+Dragon 0.053 0.169 0.109 0.094 2.483 7.092
+Earth 0.051 0.185 0.102 0.104 2.232 7.139
+VI Application
+We validate the effectiveness of our algorithm through five applications.
+VI-A Point-to-Mesh Distance Queries
+Given a mesh surface, fast querying of the closest geometric primitive (vertex, edge, or face) to a user-specified point, along with the corresponding closest point and minimum distance, represents a fundamental operation across diverse research domains [31, 32, 33, 34] including computer graphics, physical simulation, computational geometry, and computer-aided design.
+P2M [3] currently stands as the fastest point-to-mesh distance query algorithm.
+During preprocessing, P2M constructs Voronoi diagrams of model point clouds and builds interception tables for each point cloud’s Voronoi cell.
+In the query phase, it first employs a KD-tree to compute the nearest model vertex, then utilizes the interception table to obtain the nearest distance to the mesh.
+In our experiments, we replace the KD-tree component with our algorithm.
+Since P2M already requires Voronoi diagram construction, our data structure introduces minimal additional preprocessing overhead while achieving 1-5× query speedup.
+We validate these speed improvements on the ABC dataset and several standard models, comparing against BVH-based implementations from Libigl [35] and FCPW [36] libraries (with vectorization enabled), as well as the original P2M method.
+Following the experimental setup of the P2M paper, for each test model, we randomly sample a million query points within the 10× axis-aligned bounding box.
+Query time results are presented in Figure 17 and Table V.
+The modified P2M algorithm demonstrates significant efficiency gains, achieving 1-5× speedup over the original P2M query performance.
+Preprocessing times are shown in Figure S1 and Table S1 of the supplementary material, where our preprocessing overhead remains nearly identical to the original P2M implementation.
 […]
-Figure 9: Comparison about query performance on ABC dataset.
-Top: the average timing cost per query ($\mu$s) for KD Tree, R Tree and Ours.
-The y-axis is scaled logarithmically.
-Bottom: the comparison about the query cost among KD Tree, R Tree and Ours.
-Dashed line indicates where the ratio equals 1.
+Figure 17: Query time comparison for point-to-mesh distance computation on the ABC dataset.
+TABLE V: Query time comparison ($\mu$s) of different methods for point-to-mesh distance computation on standard models.
+The highest and second-fastest query times are highlighted in bold and underlined, respectively
+camel bunny armadillo dragon sponza kitten lucy
+Faces 19510 69451 99976 249882 262267 268896 525814
+Libigl 3.104 7.032 5.521 3.897 0.813 17.03 4.350
+FCPW 0.933 1.984 1.617 1.061 0.213 4.537 1.338
 […]
-Figure 10: Comparison about query performance on Thingi10K data.
-Top: the average timing cost per query ($\mu$s) for KD Tree, R Tree and Ours.
-The y-axis is scaled logarithmically.
-Bottom: the comparison about the query cost among KD Tree, R Tree and Ours.
-Dashed line indicates where the ratio equals 1.
-V-E Other Types of Data
-Uniform Point Cloud.
-To fully demonstrate the efficiency of the algorithm, we compared KD Tree, R Tree, and Ours on uniform point clouds.
-As shown in Figure 11, it can be observed that even on completely uniform point clouds, our method remains comparable in query efficiency to the fastest current implementation of KD Tree and R Tree.
 […]
-Figure 11: Average query time on the uniform point cloud with varying resolutions.
-Top: Average query time per query ($\mu$s) among KD Tree, R Tree and Ours.
-Bottom: The comparison of the average time of different methods.
-Dashed line indicates where the ratio equals 1.
-Earth City.
-Additionally, we tested the efficiency of the KD Tree, R Tree, and our method on an Earth data, distributed over a sphere, with the results shown in Figure 12.
-At this distribution, both KD Tree and R Tree exhibit a significant drop in efficiency.
-However, our method remains extremely fast, achieving up to 40x the query speed of KD Tree.
+VI-B Iterative Closest Point (ICP) Registration
+Rigid registration, which finds an optimal rigid transformation to align a source point set with a target point set, is a fundamental problem in computer vision and many other areas.
+The Iterative Closest Point (ICP) algorithm [37] is one of the most important algorithms for rigid registration.
+ICP operates iteratively, where in each iteration, every point in the source point set requires querying its nearest neighbor in the target point set, which significantly impacts the algorithm’s efficiency.
+In our experiments, we override the KD-tree implementation in PCL’s [30] ICP algorithm, substituting it with our method and nanoflann’s KD-tree implementation respectively.
+We evaluate three methods: the original ICP, ICP with our algorithm ($ICP_{ours}$), and ICP with nanoflann implementation ($ICP_{kdt}$) on 8 test scenarios from the 3DMatch dataset [38], performing registration on all adjacent frame pairs.
+The timing results are shown in Figure 18 (statistics include complete runtime including both preprocessing and querying), where the x-axis represents the sum of points in the source and target point clouds.
+Across all scenarios, our algorithm achieves faster registration in 82.4% of the cases, with speedups exceeding 10× compared to $ICP_{kdt}$ in the best cases.
+Moreover, the advantage becomes more pronounced as the number of queries increases.
 […]
-Figure 12: Average query time on the earth data with varying resolutions.
-Top: Average query time among KD tree, R tree and ours.
-Bottom: The comparison about the average time different methods.
-Dashed line indicates where the ratio equals 1.
-V-F Near and far query points.
-Additionally, we observed that the location of query points at different positions can impact the efficiency of the query.
-To further investigate this, we conducted tests on uniformly distributed point clouds, point clouds representing Earth cities, and the Dragon model, with the number of points in each point cloud controlled at 1, [^0] K.
-The results are presented in Figure 13.
-The horizontal axis represents the ratio of the length of the generation space for the query points to the length of the bounding box of the target points along any coordinate axis.
-It can be observed that even in a completely uniform point cloud distribution, where the query point generation space matches the bounding box of the target points, our query speed remains comparable to that of the KD Tree.
-In other scenarios, our method demonstrates a significant advantage in query speed.
+Figure 18: ICP registration speed comparison on 8 3DMatch test cases.
+The horizontal axis shows the sum of target and source point cloud sizes.
+Our approach achieves faster performance on 82.4% of the data.
 […]
-Figure 13: Average query time with varying sizes of the sampling box (arbitrary axis ratio, with volume ratio calculated as the cube of the axis ratio).
-Top: Uniform point cloud.
-Middle: Earth city data.
-Bottom: Dragon point cloud.
-V-G Farthest point sampling
-QuickFPS [22] is currently the fastest farthest point sampling strategy, achieving performance speedups of 230.5× over traditional CPU implementations, 43.4× over traditional GPU implementations, and 12.2× over previous state-of-the-art point cloud accelerators.
-It divides the point cloud into multiple buckets based on spatial location, and utilizes advanced techniques such as merged computation and implicit computation to accelerate the farthest point sampling process.
-Although the method proposed in this paper is not specifically designed for farthest point sampling, it also demonstrates impressive speed in this task.
-Using several traditional models as examples, with the farthest point sampling rate set at 25%, the sampling times of QuickFPS (CPU Version) and the method proposed in this paper are shown in Table IV.
-Moreover, compared to the complexity of QuickFPS, our proposed algorithm are both simpler and more intuitive.
-TABLE IV: Comparison of Sampling Times (s) between QuickFPS and Ours at a 25% Sampling Rate.
-Camel Bunny Dragon Kitten Armadillo Lucy Sponza
-Vertices 28934 72911 100313 291023 726367 1018219 1313504
-QuickFPS 0.016 0.094 0.156 1.109 2.391 2.875 3.969
-Ours 0.079 0.25 0.328 1.484 5.421 8.453 11.25
-VI Desirable Properties
-In light of its speed advantages, we believe that the current method harbors substantial potential for further advancements.
-In the following, we will delineate several pivotal aspects pertaining to its applications, aiming to elucidate the breadth and depth of its prospective utility in various domains.
-VI-A Nearest Neighbor Among the First $k$ Points
-One notable property of our algorithm is that, given any query point $q$ and integer $k$, we can efficiently identify the nearest point among $\{p_{i}\}_{i=1}^{k},k\leq n$.
-Specifically, the order of the given points is used to incrementally construct the Delaunay.
+Figure 19: Four 2D clustering datasets from [39, 40], each containing 1M points.
+VI-C Density Peak Clustering in 2D
+Clustering is the task of grouping similar objects into clusters and represents a fundamental operation in data analysis and unsupervised machine learning.
+Clustering algorithms can be used to identify different types of tissues in medical imaging [41], analyze social networks, and identify weather regimes in climatology [42].
+Among various clustering approaches, density-based clustering defines clusters as dense regions of points in the coordinate space and has received considerable attention.
+And Density peak clustering (DPC) [43] is a popular version of density-based clustering.
+A critical step in the DPC algorithm involves calculating $\delta_{i}$, which is defined as:
+$\delta_{i}=\min_{j:\rho_{j}>\rho_{i}}(d_{ij}),$ (8)
+where $\rho_{i}$ represents the density of point $p_{i}$, and $d_{ij}$ denotes the distance between points $p_{i}$ and $p_{j}$.
+Generally, computing $\{\delta_{i}\}_{i=1}^{n}$ requires calculating distances between every pair of points, resulting in $O(n^{2})$ complexity.
+Numerous algorithms have been proposed to accelerate the computation of $\delta_{i}$ through heuristic approaches [44, 45].
+One notable property of our algorithm is its ability to efficiently identify the nearest point among $\{p_{i}\}_{i=1}^{k},k\leq n$ for any query point $q$.
 As discussed in Section III, during the nearest-neighbor search process, when querying the nearest neighbor, we only need to compute $\Phi_{n}(q)$.
@@
 When a specific $k$ is provided, as shown in Algorithm 2, our query process can terminate early.
-To the best of our knowledge, no existing algorithm is capable of achieving this.
-The Density Peak Clustering algorithm [9] is a well-known method for clustering. One critical step in this algorithm is the calculation of $\sigma_{i}$, which is defined as:
-$\delta_{i}=\min_{j:\rho_{j}>\rho_{i}}(d_{ij}),$ (9)
-$\rho_{i}$ is defined as the density of point $p_{i}$, and $d_{ij}$ represents the distance between points $p_{i}$ and $p_{j}$.
-Generally, computing $\{\delta_{i}\}_{i=1}^{n}$ requires calculating the distance between every pair of points, which has a complexity of $O(n^{2})$.
-In contrast, our algorithm can efficiently solve this problem in $O(n\log n)$ time, assuming the data points are two-dimensional or three-dimensional.
-However, for higher-dimensional data, the time complexity of our algorithm increases due to the computationally expensive Delaunay triangulation process.
-VI-B Farthest Point Sampling.
-Farthest Point Sampling (FPS) is a widely used sampling algorithm because it ensures uniform point distribution, making it highly applicable across various tasks.
-For instance, FPS is used in the PointNet++ [35] framework for 3D point cloud deep learning to ensure uniform sampling and clustering of points, which contributes to building an effective receptive field.
-Although the initial purpose of our method was not specifically for farthest point sampling, it naturally lends itself to a simple and fast FPS strategy.
-Compared to existing methods, our proposed FPS strategy is more straightforward and easier to implement, offering an intuitive solution for efficient point sampling.
-VI-C Various Primitives & Distance Metrics
-Many existing nearest neighbor search algorithms are designed for specific types of primitives and distance metrics, which limits their generalizability across diverse scenarios.
-In contrast, our method, as described in Sections IIIIV-A, and IV-B, requires only the incremental construction of the Voronoi diagram for the primitives and a distance query function from any point in space to the primitives, making it straightforward to integrate into our nearest neighbor query framework.
-For instance, by supplying only the algorithm for incrementally constructing the Voronoi diagram of triangles under the Manhattan distance, along with a function to query the Manhattan distance between a spatial point and a triangle, our framework can efficiently determine the nearest primitive to a given point under the Manhattan distance.
-We believe this is a significant advantage of our algorithm.
-Of course, we acknowledge that the incremental construction of Voronoi diagrams for various primitives is quite challenging; however, as outlined below, we also provide alternative options.
-VI-D Fast Preprocessing with Compromising Query Performance
-In what follows, we aim to enhance the preprocessing efficiency at the expense of query performance.
-As elaborated in Section IV-A, during the construction of the Query Table, it is essential to identify the set $\omega_{m}$ upon inserting $p_{m}$, as illustrated in Algorithm 1.
-The set is defined by:
-$\omega_{m}=\{p_{i}~{}|~{}p_{m}\vdash\text{Cell}(p_{i};\mathcal{V}_{m-1})\}.$ (10)
-Indeed, when confronted with complex high-dimensional or intricate scenarios, computing $\omega_{m}$ becomes exceedingly challenging.
-Consequently, we opt to compute $\omega_{m}^{*}$, which merely needs to satisfy $\omega_{m}\subseteq\omega_{m}^{*}$, while still ensuring accurate nearest neighbor results.
-This approach opens up avenues for executing rapid nearest neighbor searches in high-dimensional spaces and diverse primitive and distance metric contexts.
-As an inevitable outcome, the computed set may harbor redundancies, thereby causing a marginal decrease in the speed of the nearest neighbor search.
-Moreover, even if $\omega_{m}$ does not fall within $\omega_{m}^{*}$, it could furnish valuable insights conducive to the development of approximate nearest neighbor search algorithms.
-VI-E Fast Query with Elaborated Preprocessing
-Another encouraging finding is that the number of comparisons in our query process, as shown in Table I, can be further reduced, suggesting potential for additional algorithmic acceleration.
-Specifically, as previously outlined in our algorithm, when we focus on the time point $\Phi_{k}(q)=p_{k}$ during the query process, the current approach only utilizes the information that $q\in\text{Cell}(\Phi_{k}(q);\mathcal{V}_{k})$.
-However, we can infer that $q\in\cap_{\Phi_{i}(q)=p_{i},1<i\leq k}\text{Cell}(\Phi_{i-1}(q);\mathcal{V}_{i-1})~{}\cap~{}\text{Cell}(\Phi_{k}(q);\mathcal{V}_{k})$ from the query process.
-As illustrated in Figure 14, (a) and (b) depict a simple incremental construction process, with the red dot representing the query point $q$.
-In this example, we have only utilized the information shown in (c), namely that $q\in\text{Cell}(p_{3};\mathcal{V}_{3})$.
-As shown in brown in (c).
-However, based on the query process, we can infer that $q\in\text{Cell}(p_{1};\mathcal{V}_{2})~{}\cap~{}\text{Cell}(p_{3};\mathcal{V}_{3})$, as is shown in (d).
-This enables us to design a more rigorous and precise evaluation criterion.
-Our method currently does not leverage this information during the query process, even though it can be obtained through complex preprocessing, which would, of course, involve substantial preprocessing overhead.
+To the best of our knowledge, no existing algorithm achieves this capability.
+Therefore, in 2D spaces, our method is particularly well-suited for accelerating density peak clustering algorithms.
+We simply sort the point cloud by density and use this ordering for incremental preprocessing in our algorithm.
+In our experiments, we replace the $\delta$ computation process in Ex-DPC++ [44], the fastest known density peak clustering algorithm as we known, with our method.
+We evaluate performance on four classical scenarios [39, 40], each containing 100,000 points, as shown in Figure 19, and 1000 randomly generated 2D clustering scenarios.
+Table VI and Figure 20 present the comparative results.
+Without parallelization, our clustering achieves over 10× speedup compared to the original Ex-DPC++.
+Even with parallelization enabled (while our $\delta$ computation remains serial), our method maintains 1-3× speed advantages.
+It should be noted that due to the complexity of high-dimensional Delaunay construction, our strategy-accelerated density peak clustering algorithm is currently suitable only for low-dimensional spaces (2D/3D), with high-dimensional clustering requiring further investigation.
+Beyond the speed advantages, our method offers stronger theoretical foundations compared to existing heuristic-based and parallelized algorithms.
+Additionally, our capability to query the nearest point among the first $k$ points may find applications in various practical scenarios:
+  1. 1.
+    Finding the most similar weather conditions to current patterns that occurred before a specific date for climate analysis.
+  2. 2.
+    Identifying the closest historical stock price pattern to current market conditions, restricted to periods before specific regulatory changes.
+  3. 3.
+    In mobile location services, users seek restaurants that are both “affordable and nearby.” The system needs to identify the nearest restaurant among all lower-priced establishments..
+  4. 4.
+    Finding “highly qualified and role-appropriate” benchmark employees for job candidates during recruitment processes.
+  5. 5.
+    Journal editors need to identify reviewers who are both “academically distinguished and research-field relevant” for submitted manuscripts.
 […]
-Figure 14: (a) and (b): A simple incremental insertion process, with the red dot representing the query point $q$.
-(c) Based on the query process of (a)(b), we know the information $q\in\text{Cell}(p_{3};\mathcal{V}_{3})$, as shown in brown.
-(d) In fact, the complete information we can obtain is $q\in\text{Cell}(p_{1};\mathcal{V}_{2})~{}\cap~{}\text{Cell}(p_{3};\mathcal{V}_{3})$, as shown in brown.
-However, we also recognize that utilizing all of the aforementioned information would result in significantly increased preprocessing complexity.
-To balance this, we implemented a simplified approach that compromises between the algorithm described in this paper and the information mentioned in this section.
-In brief, in addition to $q\in\text{Cell}(p_{i};\mathcal{V}_{i})$, we incorporated $q\in\text{Cell}(\Phi_{i-1};\mathcal{V}_{i-1})~{}\cap~{}\text{Cell}(p_{i};\mathcal{V}_{i})$.
-This region is used as the fundamental unit for occupancy evaluation, rather than a whole Voronoi cell.
-We were pleasantly surprised to find that, despite implementing only a simplified version, we achieved a significant reduction in the number of comparisons compared to Table I.
-As shown in Table V, $Ours^{1}_{*}$ and $Ours^{2}_{*}$ represent the number of comparisons without and with farthest point sampling, respectively, under the strategy outlined in this section.
-Remarkably, only around 100 pairwise distance comparisons are required to find the exact nearest neighbor for a query point among 1 million points.
-However, we must acknowledge that while the number of comparisons has decreased, the actual query time has increased.
-We speculate that this may be due to increased memory usage by the data structure during the query phase, which likely results in more cache misses, thereby reducing efficiency.
-Nevertheless, we believe this is a highly meaningful direction for further exploration.
-TABLE V: The number of pairwise comparisons during the query process using different construction strategies.
+Figure 20: Clustering speed comparison on 1000 synthetic clustering data examples, without (up) and with (bottom) parallelization.
+Our method demonstrates consistent performance advantages in both scenarios.
+TABLE VI: Clustering speed comparison (in seconds) on four clustering data examples, before and after OpenMP parallelization. Our method demonstrates consistent speed advantages across all examples.
+syn-2d Birch1 Birch2 Birch3
+Ours 0.715 0.432 0.577 0.796
+Ex-DPC++ 4.974 5.723 4.961 5.386
+Ours (omp) 0.559 0.298 0.407 0.610
+Ex-DPC++ (omp) 0.927 0.953 0.1008 0.960
+VI-D Point-to-Segments Distance Query
+In 2D scenarios, computing the nearest distance from a point to a collection of line segments represents a fundamental task in computer graphics, with widespread applications in path planning, navigation, and collision detection.
+Traditional approaches for nearest segment queries typically employ hierarchical bounding volume strategies, such as bounding volume hierarchies (BVH) or R-trees, to efficiently organize and traverse the line segment collection.
+Although this paper primarily focuses on computing nearest distance queries from points to point sets on 2D manifold surfaces in 3D space, our nearest neighbor search framework can be extended to different primitives, dimensions, distance metrics, and even solution spaces, provided that incremental Delaunay construction algorithms exist for the specific scenario and a distance query function from any point in the computational space to the primitives is available.
+Leveraging CGAL’s integrated 2D segment Delaunay construction algorithm [46], our nearest neighbor search approach can be adapted for point-to-disjoint-segments nearest distance queries in 2D.
+In our experiments, we randomly generate 50,000 non-intersecting line segments within the domain $[-5,5]^{2}$, with segment lengths randomly distributed in the range $[0,\delta]$.
+By controlling the parameter $\delta$, we obtain multiple test scenarios with varying geometric complexity.
+We randomly generate 1 million query points within $[-5,5]^{2}$ and compare against FCPW and R*Tree (from Boost).
+Figure 21 presents the comparative results, where the x-axis represents the percentage of $\delta$ relative to the diagonal length of the generation space ($5\sqrt{2}$), and the y-axis shows the average query time per operation.
+As observed, both R*Tree and BVH-based methods exhibit declining query efficiency as the maximum segment length increases.
+In contrast, our method maintains consistent query efficiency across scenarios with different maximum segment lengths while demonstrating superior performance compared to existing approaches.
+Preprocessing times are shown in Figure S2 of the supplementary material.
+[…]
+Figure 21: Point-to-segments distance query performance comparison with varying maximum segment lengths.
+VI-E Farthest Point Sampling as a Natural Extension
+Although our method was not specifically designed for farthest point sampling, our incremental Voronoi construction framework naturally extends to support this functionality.
+The algorithm can readily identify farthest points during the incremental construction process, making it a straightforward derivation of our core approach.
+While our naturally derived implementation is slower than existing specialized farthest point sampling algorithms [47], it demonstrates the versatility of our framework and provides a simple solution that emerges directly from our Voronoi construction methodology.
+The comparative performance results are presented in Table VII.
+TABLE VII: Farthest point sampling time comparison between QuickFPS and our method at 25% sampling rate on standard models.
 Camel Bunny Dragon Kitten Armadillo Lucy Sponza
 Vertices 28934 72911 100313 291023 726367 1018219 1313504
-$\text{Ours}^{1}$ 171.783 166.726 209.516 182.719 234.742 276.549 418.285
-$\text{Ours}^{2}$ 144.754 135.831 161.822 155.806 197.042 235.085 358.413
-$\text{Ours}^{1}_{*}$ 93.218 97.49 108.545 107.708 127.760 134.123 124.615
-$\text{Ours}^{2}_{*}$ 75.453 74.055 88.089 82.812 94.953 106.269 108.203
+QuickFPS 0.023 0.064 0.087 0.265 0.865 1.401 2.013
+Ours 0.035 0.102 0.142 0.610 2.361 3.428 4.501
 VII Limitations and Future Work
-The primary limitation of our algorithm is its high preprocessing cost, which restricts its applicability, particularly in large-scale scenarios or with high-dimensional datasets where efficiency is crucial.
+The primary limitation of our algorithm is its high preprocessing cost and the difficulty of parallelizing the preprocessing phase, which restricts its applicability, particularly in large-scale scenarios or with high-dimensional datasets where efficiency is crucial.
 Therefore, in future work, we plan to couple the construction of our algorithm more closely with Delaunay computations, or alternatively, develop an approximate Delaunay construction to achieve faster preprocessing.
-These approaches aim to reduce the preprocessing burden and improve scalability for large-scale and high-dimensional applications.
+Additionally, we will explore the potential for accelerating the preprocessing phase through parallelization strategies.
+These objectives aim to reduce the preprocessing burden and improve scalability for large-scale and high-dimensional applications.
 Additionally, our method currently only supports nearest neighbor search.
 In future work, we aim to extend the algorithm to support k-nearest neighbor queries, which would enhance its versatility.
-Moreover, addressing all the desirable properties outlined in Section VI remains a significant long-term goal for us.
+Furthermore, our approach is primarily effective in low-dimensional spaces (2D/3D), as high-dimensional Delaunay triangulation becomes computationally expensive with prohibitive preprocessing costs.
+Exploring approximate Delaunay construction techniques for high-dimensional spaces represents another promising direction that could potentially overcome this dimensional constraint while maintaining reasonable preprocessing costs.
 VIII Conclusion
@@
 Inspired by the incremental Delaunay construction process, we propose a novel nearest neighbor search algorithm specifically designed for 2D manifold point cloud data.
-Our method achieves a speedup of 1-10 times compared to the KD Tree approach.
-We validated the efficiency advantages of our approach through extensive experiments.
-The high performance of our algorithm is also supported by evidence showing that it requires only about 100 pairwise distance comparisons to identify the point closest to the query point from a set of 1 million points.
-Additionally, our algorithm demonstrates significant potential.
-For instance, it supports querying the closest point to $q$ among any of the first $k$ points.
-This capability facilitates the application of peak clustering algorithms to large-scale data scenarios.
-Moreover, we have implemented a farthest point sampling algorithm with $O(k\log n)$ complexity in 3D space.
-We believe that the algorithm presented in this paper holds significant potential for application and will make a substantial contribution to the community.
+Our method achieves 1-10× speedup compared to KD-Tree and R*-Tree approaches, with efficiency advantages validated through extensive experiments.
+Additionally, our algorithm demonstrates significant practical potential.
+We validated the effectiveness and versatility of our approach through five diverse applications: point-to-mesh distance queries, ICP registration, 2D density peak clustering, point-to-segments distance queries, and farthest point sampling.
+We believe that the algorithm presented in this paper offers substantial practical value and theoretical contributions that will benefit the broader research community.